### v1

# On quantum learning algorithms for noisy linear problems

Abstract
Quantum algorithms have shown successful results in solving noisy linear problems with quantum samples in which cryptographic hard problems are relevant. In this paper the previous results are investigated in detail, leading to new quantum and classical algorithms under the same assumptions as in the earlier works. To be specific, we present a polynomial-time quantum algorithm for solving the ring learning with errors problem with quantum samples which was deemed to be infeasible in [12], as well as polynomial-time classical algorithms that are more efficient than the corresponding quantum algorithms in solving the short integer solution problem with quantum samples and the learning with errors problem with size-reduced quantum samples.
Keywords:
Quantum algorithm Bernstein-Vazirani algorithm Learning with errors Machine learning
1 Introduction
In designing algorithms, setting underlying assumptions on what one can and cannot do is a crucial starting point. Indeed in a study of quantum algorithms, seemingly innocent assumptions sometimes become the center of controversies, for example one in a blackbox query [7, 9], in a state preparation [1], or on quantum random access memory (QRAM) [13]. Considering that the field is not in mere proof-of-concept stage, it is a necessary task to examine such assumption and evaluate its plausibility.
One of the rapidly growing fields in quantum technology is machine learning where researchers try to make the best use of physical phenomena in studying the patterns in data. In 2019, an important milestone was achieved by Grilo et al. [12] where the authors have proven that the learning with errors (LWE) problem is efficiently solvable by using Bernstein-Vazirani (BV) algorithm [5], if a certain form of a superposed data sample is available. The paper also investigated the applicability of the learning algorithm to the learning parity with noise (LPN), the ring learning with errors (RLWE), the learning with rounding, and the short integer solution (SIS) problems, although not all were successful. The algorithm was further revised in 2022 by utilizing the divide-and-conquer strategy, basically dealing with component-wise problems [19].
As clearly noted by the authors of [12], it does not directly mean the end of some lattice based cryptography, as it is not clear how such sample can be prepared. There has been an attempt to build a quantum circuit for QRAM in order to estimate resources for the sample preparation [20], but the very design requires exponentially many qubits which likely hinder the realization of the attack. On the opposite hand, there has been no mathematical proof that an efficient way of obtaining quantum samples defined in [12] is impossible. To sum up, it is too early to be optimistic (or pessimistic) on the plausibility of the learning algorithms along the line.
Nevertheless, we have noticed in some occasions specific to [12, 19], certain assumptions are highly unlikely, or the negative prospect given by the authors turns out to be feasible. This work is summarized as follows:
  • Grilo et al. have argued in their paper (Appendix B in [12]) that their LWE algorithm cannot be extended to solve RLWE problems under the error model by Bshouty and Jackson [8]. After examining the issues and problems raised by the authors, we have come up with a solution leading to a quantum learning algorithm for the RLWE problem.
  • The same authors have claimed that the SIS problem can be solved by the learning algorithm with quantum samples. Here we show our concern that a quantum sample assumption in SIS is less plausible than that in LWE. In particular we show that given the SIS sample introduced by the authors, there exists a polynomial-time classical algorithm for the SIS problem.
  • Divide-and-conquer algorithm proposed by [19] takes advantage of component-wise approach to the secret vector, suggesting size-reduced quantum samples. We question the practicality of such samples by proposing classical algorithms for the LWE problem given the same samples.
The rest of the paper is organized as follows. In Section 2, we define some notations and introduce hard problems relevant to noisy linear problems. A brief review of the Bernstein-Vazirani algorithm and how it brings an algorithm for solving the LWE problem with quantum samples are also covered. Section 3 introduces a new algorithm for solving the RLWE problem with quantum samples. An algorithm for the SIS problem with quantum samples devised by Grilo et al. [12] is reexamined in Section 4. In Section 5, we give polynomial-time classical algorithms for the LWE problem with size-reduced samples.
2 Preliminaries
This section first introduces cryptographic hard problems related with noisy linear learning, and reviews the (generalized) Bernstein-Vazirani (BV) algorithm [5] followed by the algorithm of [12] for solving the LWE problem with quantum samples. Readers are assumed to be familiar with Bra-ket notation and the elementary quantum computation.
The following notation is used throughout the paper.
  • $\bullet$
    For $q>2$, $\mathbb{Z}_{q}:=[-\frac{q}{2},\frac{q}{2})\cap\mathbb{Z}$. For $q=2$, $\mathbb{Z}_{2}:=\{0,1\}$. We also define +=$\mathbb{Z}^{+}=\{a\in\mathbb{Z}:a>0\}$.
  • $\bullet$
    An inner product of two vectors ${\bm{a}}=(a_{0},a_{1},\ldots,a_{n-1})$, ${\bm{b}}=(b_{0},b_{1},\ldots,b_{n-1})$ is defined as follows:
    ${\bm{a}},{\bm{b}}\right>\;$ $=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{n-1}b_{n-1}.$
  • $\bullet$
    Let $x\leftarrow\chi$ denote a variable $x$ taking a value according to the probability distribution $\chi$, and let $x_{0},\ldots,x_{l-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ denote that $x_{0},\ldots,x_{l-1}$ are sampled independently according to the probability distribution $\chi$. Let $\mathcal{U}(S)$ denote the uniform distribution on a set $S$. Let $\mathcal{B}_{\eta}$ denote the Bernoulli distribution with parameter $\eta\in[0,1/2)$ so that $\Pr[x=b:x\leftarrow\mathcal{B}_{\eta}]=b\!\cdot\!\eta+(1-b)\!\cdot\!(1-\eta)$ for $b\in\{0,1\}$.
  • $\bullet$
    Let us denote $\mathsf{poly}(x_{1},\ldots,x_{m})=O(x_{1}^{c_{1}}\cdots x_{m}^{c_{m}})$ for some constants $c_{1},\ldots,c_{m}$.
2.1 Hard problems
Here we describe cryptographic hard problems such as LPN, LWE, RLWE, and SIS. The hardness of these problems are guaranteed in the following sense: LPN problem can be seen as the average-case analogue of the decoding random linear codes which is NP-complete [4], and LWE, RLWE, SIS problems are known to be at least as hard as certain lattice problems in worst-case [17, 6, 15, 2, 16] when they are properly parameterized.
2.1.1 Learning parity with noise
Let ${\bm{s}}\in\{0,1\}^{n}$ be a secret vector and let $\eta\in[0,1/2)$ be a noise parameter. An LPN sample is given by
$({\bm{x}},\left<{\bm{s}},{\bm{x}}\right>+e\bmod 2)\in\{0,1\}^{n}\times\{0,1\}$
where ${\bm{x}}\leftarrow\mathcal{U}(\{0,1\}^{n})$ and $e\leftarrow\mathcal{B}_{\eta}$. The (search) LPN problem asks to find the secret vector ${\bm{s}}\in\{0,1\}^{n}$, given such LPN samples.
2.1.2 Learning with errors
The LWE problem is a generalization of the LPN problem. Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector where $q\in\mathbb{Z}^{+}$, and let $\chi_{\textup{lwe}}$ be a probability distribution on $\mathbb{Z}$, called error distribution. An LWE sample is given by
$({\bm{a}},\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$
where ${\bm{a}}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and $e\leftarrow\chi_{\textup{lwe}}$. The (search) LWE problem asks to find ${\bm{s}}$, given such LWE samples.
2.1.3 Ring learning with errors
The RLWE problem is similar to the LWE problem, but it deals with elements in the ring $\mathcal{R}:=\mathbb{Z}[X]/\langle\phi(X)\rangle$ instead of a vector in $\mathbb{Z}_{q}^{n}$, [^1] The RLWE problem is in general defined on a ring of integers of a number field, but for simplicity only polynomial rings are considered in this work. It is well-known that these two definitions are equivalent [18]. where $\phi(X)\in\mathbb{Z}[X]$ is a monic irreducible polynomial of degree $n$. Let $s(X)\in\mathcal{R}_{q}:=\mathcal{R}/q\mathcal{R}$ be a secret polynomial where $q\in\mathbb{Z}^{+}$. Let $\chi_{\textup{rlwe}}$ be a probability distribution on $\mathcal{R}$ which samples $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in\mathcal{R}$ as $a_{0},\ldots,a_{n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$. An RLWE sample is given by
$\big{(}a(X),s(X)\!\cdot\!a(X)+e(X)\bmod q\big{)}\in\mathcal{R}_{q}\times\mathcal{R}_{q}$
where $a(X)\leftarrow\mathcal{U}(\mathcal{R}_{q})$, $e(X)\leftarrow\chi_{\textup{rlwe}}$, and $\cdot$, $+$ denote multiplication and addition in $\mathcal{R}$, respectively. The (search) RLWE problem asks to find $s(X)$, given such RLWE samples.
2.1.4 Short Integer Solution
The SIS problem is defined as follows: given $n$ uniformly selected random vectors ${\bm{a}}_{1},\ldots,{\bm{a}}_{n}\in\mathbb{Z}_{q}^{m}$ for $q\in\mathbb{Z}^{+}$, find a nonzero ${\bm{v}}=(v_{1},\ldots,v_{n})\in\mathbb{Z}^{n}$ such that
$\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{0}}\bmod q\quad\textup{and}\quad\|{\bm{v}}\|\leq\beta$
where $\sqrt{n}q^{m/n}\leq\beta<q$. The inhomogeneous version of SIS problem is to find a short ${\bm{v}}$ such that $\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{z}}\bmod q$ for given ${\bm{z}}\in\mathbb{Z}_{q}^{m}$ where $m<n$.
2.2 Bernstein-Vazirani algorithm
In this subsection, first we review how the BV algorithm works on binary vectors, and then generalize it.
2.2.1 Binary case
Consider a vector ${\bm{s}}\in\{0,1\}^{n}$ and a function $f_{\bm{s}}:\{0,1\}^{n}\rightarrow\{0,1\}$, $f_{\bm{s}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod 2$. The problem to be tackled is
Given quantum oracle access to the unitary $U_{f_{\bm{s}}}:\lvert{\bm{x}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{x}}\rangle\lvert f_{\bm{s}}({\bm{x}})\rangle$, find ${\bm{s}}$.
Figure 1 (a) describes a quantum circuit for the BV algorithm [10, 5]. The initial state is $\lvert{\bm{0}}\rangle\lvert 1\rangle$, and then applying the gates as shown in Fig. 1 (a) achieves the goal. Details are not covered here, but the quantum state just before the measurement in Fig. 1 (a) reads,
$\frac{1}{2^{n}}\sum_{{\bm{y}}\in\{0,1\}^{n}}\sum_{{\bm{x}}\in\{0,1\}^{n}}(-1)^{\left<{\bm{s}}+{\bm{y}},{\bm{x}}\right>}\lvert{\bm{y}}\rangle\lvert-\rangle\enspace,$ (1)
where $\lvert-\rangle=\frac{1}{\sqrt{2}}(\lvert 0\rangle-\lvert 1\rangle)$. As can be seen from Eq. (1), when measuring $\lvert{\bm{y}}\rangle$ register, the only non-trivial amplitude occurs for ${\bm{y}}={\bm{s}}$, and there does not exist any chance of measuring a string other than ${\bm{s}}$. Assuming single qubit gates are simultaneously executable and the query is efficient, the time complexity of the algorithm is $O(1)$.
Figure 1: (a) A circuit description of BV algorithm and (b) its variant. In (b), the Hadamard gate on the bottom wire and the subsequent measurement are not necessary, but included for easier analysis.
Slightly different circuit is introduced in Fig. 1 (b). Using this circuit, the state before the measurement reads,
$\frac{1}{\sqrt{2}}\left(\lvert{\bm{0}}\rangle\lvert 0\rangle+\lvert{\bm{s}}\rangle\lvert 1\rangle\right)\enspace.$ (2)
Notice that the Hadamard gate acting on the bottom wire is redundant, but only there to make Eq. (2) simple. It can be seen from Eq. (2) that measuring the first register gives ${\bm{s}}$ with probability 1/2. The procedure can simply be repeated until we get ${\bm{s}}$.
2.2.2 Generalized BV algorithm
Instead of qubits, consider $q$-dimensional qudits. Each wire in Fig. 1 represents $q$-dimensional complex vectors and Hadamard gate is accordingly generalized [21]. Recall that generalized Hadamard operation can be understood as quantum Fourier transform on $\mathbb{Z}_{q}$[14] such as
$\mathsf{QFT}_{\mathbb{Z}_{q}}:\lvert x\rangle$ $\mapsto$ $\sum_{y\in\mathbb{Z}_{q}}e^{2\pi\imath\frac{xy}{q}}\lvert y\rangle\enspace.$
We are ready to describe the generalized BV algorithm. Notice that after the query in Fig. 1 (b), we would have $\frac{1}{\sqrt{q^{n}}}\sum_{{\bm{x}}\in\mathbb{Z}_{q}^{n}}\lvert{\bm{x}}\rangle\lvert f_{{\bm{s}}}({\bm{x}})\rangle$ where $f_{{\bm{s}}}:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{Z}_{q}$, $f_{{\bm{s}}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod q$. Instead of specifying a circuit for preparing such states, we simply define a quantum sample $\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$, and take it as an input to the BV algorithm as follows:
Algorithm 1 $\mathsf{BV}_{q,n}$
$\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$
2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
3:
4:Apply $\mathsf{QFT}_{\mathbb{Z}_{q}}^{n+1}$ on $\lvert\psi\rangle$
5:Measure all qudits to get $({\bm{x}},y)\in\mathbb{Z}_{q}^{n}\!\times\!\mathbb{Z}_{q}$
6:if $\gcd(y,q)=1$ then
7:     return $\tilde{{\bm{s}}}:=-y^{-1}{\bm{x}}\bmod q$triangleright component-wise modular division
8:end if
9:return $\bot$
One may think of the input $\lvert\psi\rangle$ as what we get by querying an oracle $\mathcal{O}_{\bm{s}}$. This algorithm will be the core subroutine in the learning algorithms.
At this point, it will be a natural question to ask if the algorithm still works for noisy input samples. The main contribution of [12] is to answer the question as briefly reviewed below.
2.3 LWE algorithm with quantum samples
2.3.1 Quantum sample
Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector and let $\mathcal{V}\subseteq\mathbb{Z}_{q}^{n}$. A quantum LWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$, is defined as follows:
$\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{{\bm{a}}\in\mathcal{V}}\lvert{\bm{a}}\rangle\lvert\left<{\bm{s}},{\bm{a}}\right>+e_{\bm{a}}\bmod q\rangle,\quad e_{\bm{a}}\leftarrow\chi_{\textup{lwe}}$ (3)
where $\chi_{\textup{lwe}}$ is a probability distribution on $\mathbb{Z}$. We assume that the support of $\chi_{\textup{lwe}}$ is $[-k,k]\cap\mathbb{Z}$ for some small positive integer $k$, i.e. $|e_{{\bm{a}}}|\leq k$. Define a quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ which outputs $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ when queried.
2.3.2 Algorithm
Given access to the quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$, one can find the secret vector ${\bm{s}}$ by using the $\mathsf{qLWE}\textup{-}\mathsf{Solver}$ described in Algorithm 2.
Algorithm 2 $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
1:oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
3:
4:$i\leftarrow 0$
5:repeat
6:     $\lvert\psi^{\textup{lwe}}_{{\bm{s}},V}\rangle\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright oracle query
7:     $\tilde{{\bm{s}}}\leftarrow\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$
8:     if  $\tilde{{\bm{s}}}=\bot$  then
9:         return $\bot$
10:     end if
11:     if  $\mathsf{Test}(\tilde{{\bm{s}}},\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}})=1$ then
12:         return $\tilde{{\bm{s}}}$
13:     end if
14:     $i\leftarrow i+1$
15:until $i\leq\ell_{1}$
16:return $\bot$
The subroutine $\mathsf{Test}$ in Algorithm 2 is designed to check if the candidate $\tilde{{\bm{s}}}$ is indeed the secret vector ${\bm{s}}$ by using additional LWE samples. Note that in Algorithm 3 the oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ can be quantum or classical.
Algorithm 3 $\mathsf{Test}$
1:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$, oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
2:$1$ (pass) or 00 (fail)
3:
4:$i\leftarrow 0$
5:repeat
6:     $({\bm{a}},b)\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright $b=\left<{\bm{a}},{\bm{s}}\right>+e_{\bm{a}}\bmod q$
7:     if $|b-\left<{\bm{a}},\tilde{{\bm{s}}}\right>\bmod q|>k$ then
8:         return 00
9:     end if
10:     $i\leftarrow i+1$
11:until $i\leq\ell_{2}$
12:return $1$
2.3.3 Performance
Detailed analysis on Algorithm 2 is not covered here. Interested readers are encouraged to refer to Section IV B in [12]. Results and parameter choices are summarized below.
Success probability of $\mathsf{BV}_{q}$
Probability that $\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$ returns the correct secret ${\bm{s}}$ is lower bounded by
$\mathrm{\Omega}\!\left(\frac{|\mathcal{V}|}{kq^{n}}\right)\enspace.$
Pass/fail probability of $\mathsf{Test}$
If $\tilde{{\bm{s}}}={\bm{s}}$, $\mathsf{Test}$ returns $1$ (pass) with certainty. If $\tilde{{\bm{s}}}\neq{\bm{s}}$, the probability that $1$ is returned is at most
$\left(\frac{2k\!+\!1}{q}\right)^{\!\ell_{2}}\enspace.$
Parameter choices for $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
By setting $\mathcal{V}=\mathbb{Z}_{q}^{n}$, $\ell_{1}=O(k\log(1/\eta))$, and $\ell_{2}=1$, we have
  • success probability: $1-\eta$
  • sample complexity: $O(k\log(1/\eta))$
  • time complexity: $O(k\log(1/\eta)\cdot\mathsf{poly}(n,\log q))$  .
In the last item, the time complexities of sample preparation (i.e. oracle query), $\mathsf{BV}_{q,n}$, and $\mathsf{Test}$ are assumed to be $O(1)$, $\mathsf{poly}(n,\log q)$, and $\mathsf{poly}(n,\log q)$, respectively.
3 An RLWE algorithm with quantum samples
In this section it will be shown that the RLWE problem with quantum samples is reduced to the LWE problem with quantum samples by efficiently transforming a RLWE sample to an LWE sample. The reduction utilizes linearity of operations in $\mathcal{R}$. To be more specific, a RLWE sample is transformed to an LWE sample in a way that coefficients of a polynomial becomes a vector. The LWE algorithm described in Section 2.3 is then applied to recover $s(X)$. Grilo et al. mentioned in their paper [12] that the LWE algorithm is not applicable to the RLWE problem due to the absence of a proper isomorphism from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$. Here we show that it can be done by employing the linearity.
3.0.1 Quantum sample
Let ${\bm{s}}(X)$ be a secret polynomial and let $\mathcal{V}\subseteq\mathcal{R}_{q}$. A quantum RLWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{rlwe}}_{{\bm{s}},\mathcal{V}}\rangle$ is defined as follows:
$\lvert\psi_{s,\mathcal{V}}^{\textup{rlwe}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{a(X)\in\mathcal{V}}\lvert a(X)\rangle\lvert s(X)\!\cdot\!a(X)+e_{a}(X)\bmod q\rangle,\quad e_{a}(X)\leftarrow\chi_{\textup{rlwe}}\enspace$
where $\chi_{\textup{rlwe}}$ is a probability distribution on $\mathcal{R}$ which samples $e_{a}(X)=\sum_{i=0}^{n-1}e_{a,i}X^{i}$ as $e_{a,0},\ldots,e_{a,n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
3.0.2 Transform from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$
Recall that the polynomial ring $\mathcal{R}$ is defined as $\mathcal{R}=\mathbb{Z}[X]/\langle\phi(X)\rangle$. Let us consider a (row) vector and a matrix constructed from a polynomial $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ as follows:
$\mathsf{V}(a)=(a_{0},a_{1},\ldots,a_{n-1})\in\mathbb{Z}^{n},\quad\mathsf{M}_{\phi}(a)=\left(\begin{array}[]{ccc}&\mathsf{V}(a)&\\ &\mathsf{V}(X\!\cdot\!a\bmod\phi)&\\ &\vdots&\\ &\mathsf{V}(X^{n-1}\!\cdot\!a\bmod\phi)&\end{array}\right)\in\mathbb{Z}^{n\times n}$
where $\cdot$ is a multiplication of polynomials. Let $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$ and define
$\mathbf{P}$ $=$ $\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ \phi_{0}&\phi_{1}&\phi_{2}&\cdots&\phi_{n-1}\end{array}\right)\in\mathbb{Z}^{n\times n}\enspace.$
We then have a useful lemma.
Lemma 1
For any $a(X),b(X)\in\mathcal{R}$, the following relations hold:
$\mathsf{V}(a\cdot b\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathsf{M}_{\phi}(b)\,,$
$\mathsf{V}(X\cdot a\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathbf{P}\,.$
Proof
Straightforwardly,
$a(X)\cdot b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}X^{i}b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}\big{(}X^{i}b(X)\bmod\phi\big{)}$
which verifies the first equation.
Next, from $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$ we have $X^{n}\equiv\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi$, and thus
$X\cdot a(X)\ =\ \sum_{i=1}^{n-1}a_{i-1}X^{i}+a_{n-1}X^{n}\ \equiv\ \sum_{i=1}^{n-1}a_{i-1}X^{i}+a_{n-1}\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi\enspace.$
Therefore,
$\mathsf{V}(X\cdot a\bmod\phi)$
$=$ $(0,a_{0},\ldots,a_{n-2})+a_{n-1}(\phi_{0},\phi_{1},\ldots,\phi_{n-1})$
$=$ $(a_{0},\ldots,a_{n-1})\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ 0&0&0&\cdots&0\end{array}\right)+\ (a_{0},\ldots,a_{n-1})\left(\!\begin{array}[]{ccccc}0&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\\ \phi_{0}&\phi_{1}&\phi_{2}&\cdots&\phi_{n-1}\end{array}\!\right)$
$=$ $\mathsf{V}(a)\,\mathbf{P}\enspace.$
Before continuing, let us introduce some notations. Let $\mathbf{M}^{T}$ denote the transpose of $\mathbf{M}\in\mathbb{R}^{l\times m}$. Let $(j\!+\!1)$-th column vector of $\mathbf{M}\in\mathbb{R}^{l\times m}$ be denoted by $\mathbf{M}_{j}\in\mathbb{R}^{l\times 1}$ so that $\mathbf{M}$ can be written as $\mathbf{M}=\left(\mathbf{M}_{0}\,|\,\cdots\,|\,\mathbf{M}_{m-1}\right)$. In particular, when $l=1$, $\mathbf{M}_{j}$ denotes the $(j\!+\!1)$-th coordinate of the row vector $\mathbf{M}$.
From Lemma 1, the matrix $\mathsf{M}_{\phi}(b)$ can be written as
$\mathsf{M}_{\phi}(b)=\left(\begin{array}[]{c}\mathsf{V}(b)\\ \mathsf{V}(b)\mathbf{P}\\ \vdots\\ \mathsf{V}(b)\mathbf{P}^{n-1}\end{array}\right),$
which shows that
$(\mathsf{M}_{\phi}(b)_{j})^{T}$ $=$ $\big{(}\mathsf{V}(b)_{j},(\mathsf{V}(b)\mathbf{P})_{j},\ldots,(\mathsf{V}(b)\mathbf{P}^{n-1})_{j}\big{)}$ (8)
$=$ ($\mathsf{V}(b),((\mathbf{I}_{n})_{j})^{T}\right>,\left<\mathsf{V}(b),(\mathbf{P}_{j})^{T}\right>\ldots,\left<\mathsf{V}(b),((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\,.$
3.0.3 RLWE to LWE
Let ${\bm{a}}=\mathsf{V}(a)$ for $a\in\mathcal{R}$. Define a unitary operation,
$U_{\phi,j}:\lvert{\bm{a}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle,\quad j\in\{0,\ldots,n-1\}$
where ${\bm{a}}_{j}=(\mathsf{M}_{\phi}(a)_{j})^{T}=\big{(}\!\left<{\bm{a}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{a}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{a}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$. We see that the operation is efficiently computable as any power of $\mathbf{P}$ can be classically pre-computed by knowing the fixed polynomial $\phi(X)$.
Now let ${\bm{s}}=\mathsf{V}(s)$ and ${\bm{e}}_{a}=(e_{a,0},\ldots,e_{a,n-1})=\mathsf{V}(e_{a})$. By Lemma 1, Eq. (8), and the linearity of $\mathsf{V}$, we have
$\mathsf{V}(s\cdot a+e_{a}\bmod\phi)$ $=$ $\mathsf{V}(s)\cdot\mathsf{M}_{\phi}(a)+\mathsf{V}(e_{a})$
[…]
For a given quantum RLWE sample, we apply the following transformations:
$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}s(X)\!\cdot\!a(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
$\xmapsto{U_{\phi,j}}$ $\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle\lvert\ldots,\,\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q,\,\ldots\rangle\enspace.$
Ignoring or uncomputing irrelevant information, we have a quantum LWE sample with secret ${\bm{s}}$:
$\lvert{\bm{a}}_{j}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q\rangle\enspace.$
Alternatively, interchanging the role of $a(X)$ and $s(X)$ in the transformation, we have a quantum LWE sample with secret ${\bm{s}}_{j}$:
$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}a(X)\!\cdot\!s(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{a}},{\bm{s}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{a}},{\bm{s}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
where ${\bm{s}}_{j}:=(\mathsf{M}_{\phi}({\bm{s}})_{j})^{T}=\big{(}\!\left<{\bm{s}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{s}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{s}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$.
4 On SIS algorithm
In [12, Appendix B], a quantum sample for SIS is defined by,
$\lvert\Psi^{\textup{sis}}\rangle=\frac{1}{\sqrt{q^{nm}}}\sum_{\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}}\lvert\mathbf{A}\rangle\lvert\mathbf{A}{{\bm{v}}}\bmod q\rangle$ (9)
where $\mathbf{A}=[\,{\bm{a}}_{0}\,|\,\cdots\,|\,{\bm{a}}_{n-1}\,]$, ${\bm{a}}_{i}\in\mathbb{Z}_{q}^{m\times 1}$, and ${\bm{v}}\in\mathbb{Z}^{n}$ is a short secret vector such that $0<\|{\bm{v}}\|\leq\beta$. The authors have developed a polynomial time quantum algorithm for finding ${\bm{v}}$ by using $O(2n\beta)$ quantum samples.
The detailed description of the algorithm is not covered here. Instead, let us focus on the problem itself. From the definition of the SIS problem in Section 2.1.4, we are given an under-determined system of equations ($m<n$). Hardness of the SIS problem comes from the fact that among the exponentially many solutions ($q^{n-m}$), there exists no known way to choose a short one within polynomial time. Having ability to prepare the quantum sample multiple times for the same ${\bm{v}}$ thus already means their is a way to choose the short solution out of many solutions. On the other hand, the hardness of LWE comes from difficulty of solving noisy linear systems. Distinct LWE algorithms may require different number of noisy equations, but accessing more than one noisy equation is not implausible.
Assuming the SIS quantum sample in Eq. (9) is obtainable, a classical algorithm can be designed. By classical, we mean no quantum process is required except measurements. Let $t=O(n/m)$, and for given $t$ samples just measure the state to have
[…]
Notice that the total number of linear equations we have is $mt$, which is greater than $n$. With high probability this system of equations has the unique solution ${\bm{v}}$ in which Gaussian elimination can be applied to solve.
5 On the Divide-and-Conquer algorithm
Song et al. have published a paper in 2022 which improves the result of [12] assuming a resized sample [19]. In a nutshell, the improved algorithm and one in Section 2.3 work similarly except the samples. Therefore, here we drop the details on the algorithm and jump right into the discussion on Eq. (9) in [19], which introduces the resized sample. Recall first that an LWE sample is $({\bm{a}},b=\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$, $|e|\leq k$.
5.1 Resized sample assumption
Let $\mathcal{V}_{j}\subseteq\mathbb{Z}_{q}$, ${\bm{s}}=(s_{0},\ldots,s_{n-1})\in\mathbb{Z}_{q}^{n}$, and assume the following sample is obtainable,
$\lvert\psi_{\mathcal{V}_{j}}\rangle=\frac{1}{\sqrt{|\mathcal{V}_{j}|}}\sum_{a_{j}^{\prime}\in\mathcal{V}_{j}}\lvert a_{j}^{\prime}\rangle\lvert a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q\rangle,\quad|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)\ll q$ (10)
where $e_{j}^{\prime}$ can be considered as an error from some distribution originated from the original LWE sample.
Apparently some more assumptions are required (including QRAM) for the above statement to hold, but let us take all of them positively and Eq. (10) is indeed prepared. [^2] To us, the description of how this kind of sample is obtained in [19] is not transparent. In the following subsections, two classical algorithms for LWE problem is developed assuming the resized sample is given.
5.2 Search on the Error
Given a resized sample, a naive classical strategy is to exhaustively search $s_{j}$ with $O(qn)$ complexity (in which the scenario is indeed discussed in [19]), but more sophisticated search would be to examine the error.
The (measured) resized sample of the $j$th component reads $b_{j}^{\prime}=a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$, $|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)$. Let us consider the following procedure:
  1. 1.
    Set a value of $e_{j}^{\prime}$ such that $|e_{j}^{\prime}|\leq k^{\prime}$.
  2. 2.
    For a measured sample, compute $\tilde{s}_{j}:=(a_{j}^{\prime})^{-1}(b_{j}^{\prime}-e_{j}^{\prime})\bmod q$.
  3. 3.
    For $1\leq i\leq M$ for some integer $M$, test if $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$, where $b_{j,i}^{\prime}$ and $a_{j,i}^{\prime}$ are $i$th sample for the test on the $j$th component.
  4. 4.
    If it passes the test, return $\tilde{s}_{j}$. Otherwise, set a different value for $e_{j}^{\prime}$ and repeat.
Let us estimate the complexity. If $\tilde{s}_{j}=s_{j}$, then $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is always true and it will pass. The probability that a candidate $e_{j}^{\prime}$ leads to the correct $\tilde{s}_{j}$ is thus at least $\Omega(1/k^{\prime})$. On the other hand if $\tilde{s}_{j}\neq s_{j}$, then the probability that $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is true is $(2k^{\prime}+1)/q$, thus $\tilde{s}_{j}\neq s_{j}$ will pass with the probability $((2k^{\prime}+1)/q)^{M}$.
Compare it with the algorithm in [19]. In [19], the probability that the BV subroutine outputs $\tilde{s}_{j}=s_{j}$ is $\Omega(1/k^{\prime})$ (Eq. (17) in [19] with $|v_{j}|\approx q$). The probability that $M$-trial test accepts an incorrect answer is at most $((2k^{\prime}+1)/q)^{M}$ as in Eq. (12) in [19]. It is not difficult to see that the above classical procedure is at least as efficient as the algorithm proposed in [19].
5.3 Lattice reduction
Given two (measured) resized samples $(a_{j,1}^{\prime},b_{j,1}^{\prime}=a_{j,1}^{\prime}s_{j}+e_{j,1}^{\prime}\bmod q)$, $(a_{j,2}^{\prime},b_{j,2}^{\prime}=a_{j,2}^{\prime}s_{j}+e_{j,2}^{\prime}\bmod q)$, let us consider the following lattice:
$\mathcal{L}$ $=$ $\big{\{}{\bm{v}}\in\mathbb{Z}^{2}:{\bm{v}}\equiv x{\bm{a}}\bmod q\textup{ for some }x\in\mathbb{Z}\big{\}}$
where ${\bm{a}}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$ and ‘$\bmod~{}q$’ applies to each component. Assume that $a_{j,1}^{\prime}$ is invertible in $\mathbb{Z}_{q}$, then $(1,(a_{j,1}^{\prime})^{-1}a_{j,2}\bmod q),(0,q)$ is a basis of $\mathcal{L}$ and so $\dim(\mathcal{L})=2$, $\det(\mathcal{L})=q$.
The purpose of introducing $\mathcal{L}$ is to reduce finding $s_{j}$ to the closest vector problem (CVP). Notice that if ${\bm{e}}:=(e_{j,1}^{\prime},e_{j,2}^{\prime})={\bm{0}}$, then ${\bm{b}}:=(b_{j,1}^{\prime},b_{j,2}^{\prime})\in\mathcal{L}$. For nonzero small $\|{\bm{e}}\|$, we thus expect that ${\bm{b}}\equiv s_{j}{\bm{a}}+{\bm{e}}\bmod q$ is close to a vector $s_{j}{\bm{a}}\bmod q\in\mathcal{L}$. Therefore the task is to solve CVP in $\mathcal{L}$ for target ${\bm{b}}$.
Generally the CVP is taken to be a hard problem, but in two dimension there exists an efficient solution: Lagrange-Gauss reduction is first applied to find a good basis of $\mathcal{L}$, and then Babai’s nearest plane algorithm is used to find the closest vector in $\mathcal{L}$ to ${\bm{b}}$ [3].
By applying the Lagrange-Gauss reduction algorithm, one can compute basis vectors ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathcal{L}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(\mathcal{L})$ in time $O(\log^{3}q)$ [11, Chapter 17], where $\lambda_{i}(\mathcal{L})$ is the $i$-th successive minimum of $\mathcal{L}$. For given basis vectors $\mathbf{b}_{1},\ldots,\mathbf{b}_{\dim(\mathcal{L})}$ and a target ${\bm{u}}$, if there is a vector ${\bm{v}}\in\mathcal{L}$ satisfying $\|{\bm{v}}-{\bm{u}}\|\leq\frac{1}{2}\min_{i\in\{1,\ldots,\dim(\mathcal{L})\}}\{\|\mathbf{b}_{i}^{*}\|\}$, then the Babai’s nearest plane algorithm always outputs the vector ${\bm{v}}$ in time $O(\log^{2}q)$ [11, Chapter 18], where $\mathbf{b}_{i}^{*}$ is the $i$-th Gram-Schmidt orthogonalized vector. Assuming the Gaussian heuristic [11, Chapter 16] holds, $\|{\bm{g}}_{1}\|\approx\sqrt{\det(\mathcal{L})}=\sqrt{q}$. Since ${\bm{g}}_{1}={\bm{g}}_{1}^{*}$ and $\|{\bm{g}}_{1}^{*}\|\|{\bm{g}}_{2}^{*}\|=\det(\mathcal{L})$, we have $\|{\bm{g}}_{2}^{*}\|\approx\sqrt{q}$. Therefore Babai’s algorithm results in a vector ${\bm{v}}\in\mathcal{L}$ such that $\|{\bm{v}}-{\bm{b}}\|\leq\frac{1}{2}\sqrt{q}$. If $\|{\bm{e}}\|\leq\frac{1}{2}\sqrt{q}$, then one would expect ${\bm{v}}\equiv s_{j}{\bm{a}}\bmod q$. Indeed, since $|e_{j,i}^{\prime}|\leq k^{\prime}$ (i.e., $\|{\bm{e}}\|\leq\sqrt{2}k^{\prime}$), a condition for the above strategy to work is
$8k^{\prime 2}\leq q\enspace.$ (11)
Equation (6) in [19] assumes $k^{\prime}\ll q$, thus Eq. (11) likely hold.

### v2

@@
-On quantum learning algorithms for noisy linear problems
+Assessing the feasibility of quantum learning algorithms for noisy linear problems
 Abstract
-Quantum algorithms have shown successful results in solving noisy linear problems with quantum samples in which cryptographic hard problems are relevant.
-In this paper the previous results are investigated in detail, leading to new quantum and classical algorithms under the same assumptions as in the earlier works.
-To be specific, we present a polynomial-time quantum algorithm for solving the ring learning with errors problem with quantum samples which was deemed to be infeasible in [12], as well as polynomial-time classical algorithms that are more efficient than the corresponding quantum algorithms in solving the short integer solution problem with quantum samples and the learning with errors problem with size-reduced quantum samples.
+Quantum algorithms for solving noisy linear problems are reexamined, under the same assumptions taken from the existing literature.
+The findings of this work include on the one hand extended applicability of the quantum Fourier transform to the ring learning with errors problem which has been left open by Grilo et al., who first devised a polynomial-time quantum algorithm for solving noisy linear problems with quantum samples.
+On the other hand, this paper also shows there exist efficient classical algorithms for short integer solution and size-reduced learning with errors problems if the quantum samples used by the previous studies are given.
 keywords:
-Quantum algorithm Bernstein-Vazirani algorithm Learning with errors Machine learning
-1 Introduction
+Quantum algorithm, Quantum Fourier Transform, Learning with errors, Machine learning, Bernstein-Vazirani algorithm
+Introduction
 In designing algorithms, setting underlying assumptions on what one can and cannot do is a crucial starting point.
-Indeed in a study of quantum algorithms, seemingly innocent assumptions sometimes become the center of controversies, for example one in a blackbox query [7, 9], in a state preparation [1], or on quantum random access memory (QRAM) [13].
-Considering that the field is not in mere proof-of-concept stage, it is a necessary task to examine such assumption and evaluate its plausibility.
+Indeed in a study of quantum algorithms, seemingly innocent assumptions sometimes become the center of controversies, for example one in a blackbox query [1, 2], in a state preparation [3], or on quantum random access memory [4].
+Considering that the field is not in the mere proof-of-concept stage, it is a necessary task to examine such an assumption and evaluate its plausibility.
 One of the rapidly growing fields in quantum technology is machine learning where researchers try to make the best use of physical phenomena in studying the patterns in data.
-In 2019, an important milestone was achieved by Grilo et al.
-[12] where the authors have proven that the learning with errors (LWE) problem is efficiently solvable by using Bernstein-Vazirani (BV) algorithm [5], if a certain form of a superposed data sample is available.
-The paper also investigated the applicability of the learning algorithm to the learning parity with noise (LPN), the ring learning with errors (RLWE), the learning with rounding, and the short integer solution (SIS) problems, although not all were successful.
+Following the pioneering work on the quantum approach to learning parity with noise (LPN) problem by Cross et al.
+[5], an important milestone was achieved in 2019 by Grilo et al.
+[6] where the authors have proven that learning with errors (LWE) problem is efficiently solvable by using Bernstein-Vazirani (BV) algorithm [7], if a certain form of a superposed data sample is available.
+The paper also investigated the applicability of the learning algorithm to LPN, ring learning with errors (RLWE), learning with rounding, and short integer solution (SIS) problems, although not all were successful.
 The algorithm was further revised in 2022 by utilizing the divide-and-conquer strategy, basically dealing with component-wise problems [8].
-As clearly noted by the authors of [12], it does not directly mean the end of some lattice based cryptography, as it is not clear how such sample can be prepared.
-There has been an attempt to build a quantum circuit for QRAM in order to estimate resources for the sample preparation [20], but the very design requires exponentially many qubits which likely hinder the realization of the attack.
-On the opposite hand, there has been no mathematical proof that an efficient way of obtaining quantum samples defined in [12] is impossible.
+As clearly noted by the authors of the paper [6], it does not directly mean the end of some lattice based cryptography, as it is not clear how such a sample can be prepared.
+At the time of writing, there has been no known way to efficiently prepare such samples, but on the other hand, there has been no mathematical proof that an efficient way is nonexistent.
 To sum up, it is too early to be optimistic (or pessimistic) on the plausibility of the learning algorithms along the line.
-Nevertheless, we have noticed in some occasions specific to [12, 19], certain assumptions are highly unlikely, or the negative prospect given by the authors turns out to be feasible.
+Nevertheless, we have noticed in some occasions specific to two papers [6, 8], certain assumptions allow polynomial-time classical algorithms questioning their soundness, or the negative prospect given by the authors turns out to be feasible.
 This work is summarized as follows:
-  • Grilo et al.
-have argued in their paper (Appendix B in [12]) that their LWE algorithm cannot be extended to solve RLWE problems under the error model by Bshouty and Jackson [8].
-After examining the issues and problems raised by the authors, we have come up with a solution leading to a quantum learning algorithm for the RLWE problem.
-  • The same authors have claimed that the SIS problem can be solved by the learning algorithm with quantum samples.
-Here we show our concern that a quantum sample assumption in SIS is less plausible than that in LWE.
-In particular we show that given the SIS sample introduced by the authors, there exists a polynomial-time classical algorithm for the SIS problem.
-  • Divide-and-conquer algorithm proposed by [19] takes advantage of component-wise approach to the secret vector, suggesting size-reduced quantum samples.
-We question the practicality of such samples by proposing classical algorithms for the LWE problem given the same samples.
-The rest of the paper is organized as follows.
-In Section 2, we define some notations and introduce hard problems relevant to noisy linear problems.
-A brief review of the Bernstein-Vazirani algorithm and how it brings an algorithm for solving the LWE problem with quantum samples are also covered.
-Section 3 introduces a new algorithm for solving the RLWE problem with quantum samples.
-An algorithm for the SIS problem with quantum samples devised by Grilo et al.
-[12] is reexamined in Section 4.
-In Section 5, we give polynomial-time classical algorithms for the LWE problem with size-reduced samples.
-2 Preliminaries
-This section first introduces cryptographic hard problems related with noisy linear learning, and reviews the (generalized) Bernstein-Vazirani (BV) algorithm [5] followed by the algorithm of [12] for solving the LWE problem with quantum samples.
+  • The LWE algorithm with quantum samples is extended to solve RLWE problem, which has been deemed to be infeasible under the error model by Bshouty and Jackson [6, 9].
+  • It is shown that the quantum sample assumption in the SIS algorithm is stronger than that in LWE in the sense that a polynomial-time classical algorithm can be devised given the same sample introduced by the previous work.
+  • Divide-and-conquer algorithm takes advantage of the component-wise approach to the secret vector, assuming size-reduced quantum samples [8].
+We develop classical algorithms for the LWE problem given the same samples.
+In the following, background information is covered prioritizing introduction to the LWE algorithm with quantum samples. New findings are then followed.
+Preliminaries
+This section first introduces cryptographic hard problems related with noisy linear learning, and reviews the (generalized) Bernstein-Vazirani (BV) algorithm [7] for solving the LWE problem with quantum samples.
 Readers are assumed to be familiar with Bra-ket notation and the elementary quantum computation.
 The following notation is used throughout the paper.
-  • $\bullet$
   • For $q>2$, $\mathbb{Z}_{q}:=[-\frac{q}{2},\frac{q}{2})\cap\mathbb{Z}$. For $q=2$, $\mathbb{Z}_{2}:=\{0,1\}$. We also define +=$\mathbb{Z}^{+}=\{a\in\mathbb{Z}:a>0\}$.
-  • $\bullet$
   • An inner product of two vectors ${\bm{a}}=(a_{0},a_{1},\ldots,a_{n-1})$, ${\bm{b}}=(b_{0},b_{1},\ldots,b_{n-1})$ is defined as follows:
-    ${\bm{a}},{\bm{b}}\right>\;$ $=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{n-1}b_{n-1}.$
-  • $\bullet$
+    ${\bm{a}},{\bm{b}}\right>=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{n-1}b_{n-1}\enspace.$
   • Let $x\leftarrow\chi$ denote a variable $x$ taking a value according to the probability distribution $\chi$, and let $x_{0},\ldots,x_{l-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ denote that $x_{0},\ldots,x_{l-1}$ are sampled independently according to the probability distribution $\chi$.
@@
 Let $\mathcal{B}_{\eta}$ denote the Bernoulli distribution with parameter $\eta\in[0,1/2)$ so that $\Pr[x=b:x\leftarrow\mathcal{B}_{\eta}]=b\!\cdot\!\eta+(1-b)\!\cdot\!(1-\eta)$ for $b\in\{0,1\}$.
-  • $\bullet$
-    Let us denote $\mathsf{poly}(x_{1},\ldots,x_{m})=O(x_{1}^{c_{1}}\cdots x_{m}^{c_{m}})$ for some constants $c_{1},\ldots,c_{m}$.
-2.1 Hard problems
+  • Let $\mathsf{poly}(x_{1},\ldots,x_{m})=O(x_{1}^{c_{1}}\cdots x_{m}^{c_{m}})$ for some constants $c_{1},\ldots,c_{m}$.
+Hard problems
 Here we describe cryptographic hard problems such as LPN, LWE, RLWE, and SIS.
 The hardness of these problems are guaranteed in the following sense: LPN problem can be seen as the average-case analogue of the decoding random linear codes which is NP-complete [10], and LWE, RLWE, SIS problems are known to be at least as hard as certain lattice problems in worst-case [11, 12, 13, 14, 15] when they are properly parameterized.
-2.1.1 Learning parity with noise
-Let ${\bm{s}}\in\{0,1\}^{n}$ be a secret vector and let $\eta\in[0,1/2)$ be a noise parameter. An LPN sample is given by
-$({\bm{x}},\left<{\bm{s}},{\bm{x}}\right>+e\bmod 2)\in\{0,1\}^{n}\times\{0,1\}$
+Learning parity with noise. Let ${\bm{s}}\in\{0,1\}^{n}$ be a secret vector and let $\eta\in[0,1/2)$ be a noise parameter. An LPN sample is given by
+$({\bm{x}},\left<{\bm{s}},{\bm{x}}\right>+e\bmod 2)\in\{0,1\}^{n}\times\{0,1\}\enspace,$
 where ${\bm{x}}\leftarrow\mathcal{U}(\{0,1\}^{n})$ and $e\leftarrow\mathcal{B}_{\eta}$. The (search) LPN problem asks to find the secret vector ${\bm{s}}\in\{0,1\}^{n}$, given such LPN samples.
-2.1.2 Learning with errors
+Learning with errors.
 The LWE problem is a generalization of the LPN problem.
-Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector where $q\in\mathbb{Z}^{+}$, and let $\chi_{\textup{lwe}}$ be a probability distribution on $\mathbb{Z}$, called error distribution.
+Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector where $q\in\mathbb{Z}^{+}$, and let $\chi_{\textup{\tiny lwe}}$ be a probability distribution on $\mathbb{Z}$, called error distribution.
 An LWE sample is given by
-$({\bm{a}},\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$
-where ${\bm{a}}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and $e\leftarrow\chi_{\textup{lwe}}$. The (search) LWE problem asks to find ${\bm{s}}$, given such LWE samples.
-2.1.3 Ring learning with errors
-The RLWE problem is similar to the LWE problem, but it deals with elements in the ring $\mathcal{R}:=\mathbb{Z}[X]/\langle\phi(X)\rangle$ instead of a vector in $\mathbb{Z}_{q}^{n}$, [^1] The RLWE problem is in general defined on a ring of integers of a number field, but for simplicity only polynomial rings are considered in this work.
+$({\bm{a}},\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}\enspace,$
+where ${\bm{a}}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and $e\leftarrow\chi_{\textup{\tiny lwe}}$. The (search) LWE problem asks to find ${\bm{s}}$, given such LWE samples.
+Ring learning with errors.
+The RLWE problem is in general defined on a ring of integers of a number field, but for simplicity only polynomial rings are considered in this work.
 It is well-known that these two definitions are equivalent [16].
-where $\phi(X)\in\mathbb{Z}[X]$ is a monic irreducible polynomial of degree $n$.
+The RLWE problem is similar to the LWE problem, but it deals with elements in the ring $\mathcal{R}:=\mathbb{Z}[X]/\langle\phi(X)\rangle$ instead of a vector in $\mathbb{Z}_{q}^{n}$, where $\phi(X)\in\mathbb{Z}[X]$ is a monic irreducible polynomial of degree $n$.
 Let $s(X)\in\mathcal{R}_{q}:=\mathcal{R}/q\mathcal{R}$ be a secret polynomial where $q\in\mathbb{Z}^{+}$.
-Let $\chi_{\textup{rlwe}}$ be a probability distribution on $\mathcal{R}$ which samples $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in\mathcal{R}$ as $a_{0},\ldots,a_{n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
+Let $\chi_{\textup{\tiny rlwe}}$ be a probability distribution on $\mathcal{R}$ which samples $e(X)=\sum_{i=0}^{n-1}e_{i}X^{i}\in\mathcal{R}$ as $e_{0},\ldots,e_{n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
 An RLWE sample is given by
-$\big{(}a(X),s(X)\!\cdot\!a(X)+e(X)\bmod q\big{)}\in\mathcal{R}_{q}\times\mathcal{R}_{q}$
-where $a(X)\leftarrow\mathcal{U}(\mathcal{R}_{q})$, $e(X)\leftarrow\chi_{\textup{rlwe}}$, and $\cdot$, $+$ denote multiplication and addition in $\mathcal{R}$, respectively.
+$\big{(}a(X),s(X)\!\cdot\!a(X)+e(X)\bmod q\big{)}\in\mathcal{R}_{q}\times\mathcal{R}_{q}\enspace,$
+where $a(X)\leftarrow\mathcal{U}(\mathcal{R}_{q})$, $e(X)\leftarrow\chi_{\textup{\tiny rlwe}}$, and $\cdot$, $+$ denote multiplication and addition in $\mathcal{R}$, respectively.
 The (search) RLWE problem asks to find $s(X)$, given such RLWE samples.
-2.1.4 Short Integer Solution
+Short integer solution.
 The SIS problem is defined as follows: given $n$ uniformly selected random vectors ${\bm{a}}_{1},\ldots,{\bm{a}}_{n}\in\mathbb{Z}_{q}^{m}$ for $q\in\mathbb{Z}^{+}$, find a nonzero ${\bm{v}}=(v_{1},\ldots,v_{n})\in\mathbb{Z}^{n}$ such that
-$\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{0}}\bmod q\quad\textup{and}\quad\|{\bm{v}}\|\leq\beta$
+$\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{0}}\bmod q\quad\textup{and}\quad\|{\bm{v}}\|\leq\beta\enspace,$
 where $\sqrt{n}q^{m/n}\leq\beta<q$.
 The inhomogeneous version of SIS problem is to find a short ${\bm{v}}$ such that $\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{z}}\bmod q$ for given ${\bm{z}}\in\mathbb{Z}_{q}^{m}$ where $m<n$.
-2.2 Bernstein-Vazirani algorithm
-In this subsection, first we review how the BV algorithm works on binary vectors, and then generalize it.
-2.2.1 Binary case
-Consider a vector ${\bm{s}}\in\{0,1\}^{n}$ and a function $f_{\bm{s}}:\{0,1\}^{n}\rightarrow\{0,1\}$, $f_{\bm{s}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod 2$. The problem to be tackled is
-Given quantum oracle access to the unitary $U_{f_{\bm{s}}}:\lvert{\bm{x}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{x}}\rangle\lvert f_{\bm{s}}({\bm{x}})\rangle$, find ${\bm{s}}$.
+Bernstein-Vazirani algorithm
+Figure 1: (a) A circuit description of BV algorithm and (b) its variant
+Binary case.
+Consider a vector ${\bm{s}}\in\{0,1\}^{n}$ and a function $f_{\bm{s}}:\{0,1\}^{n}\rightarrow\{0,1\}$, $f_{\bm{s}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod 2$.
+The problem is, for given quantum oracle access to the unitary $U_{f_{\bm{s}}}:\left|{\bm{x}}\,\right\rangle\left|0\,\right\rangle\mapsto\left|{\bm{x}}\,\right\rangle\left|f_{\bm{s}}({\bm{x}})\,\right\rangle$, to find ${\bm{s}}$.
 Figure 1 (a) describes a quantum circuit for the BV algorithm [17, 7].
-The initial state is $\lvert{\bm{0}}\rangle\lvert 1\rangle$, and then applying the gates as shown in Fig.
-1 (a) achieves the goal.
-Details are not covered here, but the quantum state just before the measurement in Fig.
-1 (a) reads,
-$\frac{1}{2^{n}}\sum_{{\bm{y}}\in\{0,1\}^{n}}\sum_{{\bm{x}}\in\{0,1\}^{n}}(-1)^{\left<{\bm{s}}+{\bm{y}},{\bm{x}}\right>}\lvert{\bm{y}}\rangle\lvert-\rangle\enspace,$ (1)
-where $\lvert-\rangle=\frac{1}{\sqrt{2}}(\lvert 0\rangle-\lvert 1\rangle)$.
-As can be seen from Eq.
-(1), when measuring $\lvert{\bm{y}}\rangle$ register, the only non-trivial amplitude occurs for ${\bm{y}}={\bm{s}}$, and there does not exist any chance of measuring a string other than ${\bm{s}}$.
+The initial state is $\left|{\bm{0}}\,\right\rangle\left|1\,\right\rangle$, and then applying the gates as shown in Figure 1 (a) achieves the goal.
+Details are not covered here, but the quantum state just before the measurement in Figure 1 (a) reads,
+$\frac{1}{2^{n}}\sum_{{\bm{y}}\in\{0,1\}^{n}}\sum_{{\bm{x}}\in\{0,1\}^{n}}(-1)^{\left<{\bm{s}}+{\bm{y}},{\bm{x}}\right>}\left|{\bm{y}}\,\right\rangle\left|-\,\right\rangle\enspace,$ (1)
+where $\left|-\,\right\rangle=\frac{1}{\sqrt{2}}(\left|0\,\right\rangle-\left|1\,\right\rangle)$.
+As can be seen from Equation (1), when measuring $\left|{\bm{y}}\,\right\rangle$ register, the only non-trivial amplitude occurs for ${\bm{y}}={\bm{s}}$, and there does not exist any chance of measuring a string other than ${\bm{s}}$.
 Assuming single qubit gates are simultaneously executable and the query is efficient, the time complexity of the algorithm is $O(1)$.
-Figure 1: (a) A circuit description of BV algorithm and (b) its variant.
-In (b), the Hadamard gate on the bottom wire and the subsequent measurement are not necessary, but included for easier analysis.
-Slightly different circuit is introduced in Fig. 1 (b). Using this circuit, the state before the measurement reads,
-$\frac{1}{\sqrt{2}}\left(\lvert{\bm{0}}\rangle\lvert 0\rangle+\lvert{\bm{s}}\rangle\lvert 1\rangle\right)\enspace.$ (2)
-Notice that the Hadamard gate acting on the bottom wire is redundant, but only there to make Eq.
-(2) simple.
-It can be seen from Eq.
-(2) that measuring the first register gives ${\bm{s}}$ with probability 1/2.
+Slightly different circuit is introduced in Figure 1 (b). Using this circuit, the state before the measurement reads,
+$\frac{1}{\sqrt{2}}\left(\left|{\bm{0}}\,\right\rangle\left|0\,\right\rangle+\left|{\bm{s}}\,\right\rangle\left|1\,\right\rangle\right)\enspace.$ (2)
+Notice that the Hadamard gate acting on the bottom wire is redundant, but only there to make Equation (2) simple.
+It can be seen from Equation (2) that measuring the first register gives ${\bm{s}}$ with probability 1/2.
 The procedure can simply be repeated until we get ${\bm{s}}$.
-2.2.2 Generalized BV algorithm
+Generalized BV algorithm.
 Instead of qubits, consider $q$-dimensional qudits.
-Each wire in Fig.
-1 represents $q$-dimensional complex vectors and Hadamard gate is accordingly generalized [21].
+Each wire in Figure 1 represents $q$-dimensional complex vectors and Hadamard gate is accordingly generalized [18].
 Recall that generalized Hadamard operation can be understood as quantum Fourier transform on $\mathbb{Z}_{q}$ [19] such as
-$\mathsf{QFT}_{\mathbb{Z}_{q}}:\lvert x\rangle$ $\mapsto$ $\sum_{y\in\mathbb{Z}_{q}}e^{2\pi\imath\frac{xy}{q}}\lvert y\rangle\enspace.$
+$\mathsf{QFT}_{\mathbb{Z}_{q}}:\left|x\,\right\rangle\mapsto\sum_{y\in\mathbb{Z}_{q}}e^{2\pi\imath\frac{xy}{q}}\left|y\,\right\rangle\enspace.$
 We are ready to describe the generalized BV algorithm.
-Notice that after the query in Fig.
-1 (b), we would have $\frac{1}{\sqrt{q^{n}}}\sum_{{\bm{x}}\in\mathbb{Z}_{q}^{n}}\lvert{\bm{x}}\rangle\lvert f_{{\bm{s}}}({\bm{x}})\rangle$ where $f_{{\bm{s}}}:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{Z}_{q}$, $f_{{\bm{s}}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod q$.
-Instead of specifying a circuit for preparing such states, we simply define a quantum sample $\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$, and take it as an input to the BV algorithm as follows:
+Notice that after the query in Figure 1 (b), we would have $\frac{1}{\sqrt{q^{n}}}\sum_{{\bm{x}}\in\mathbb{Z}_{q}^{n}}\left|{\bm{x}}\,\right\rangle\left|f_{{\bm{s}}}({\bm{x}})\,\right\rangle$ where $f_{{\bm{s}}}:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{Z}_{q}$, $f_{{\bm{s}}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod q$.
+Instead of specifying a circuit for preparing such states, we simply define a quantum sample $\left|\psi\,\right\rangle\in(\mathbb{C}^{q})^{n+1}$, and take it as an input to the BV algorithm as follows:
 Algorithm 1 $\mathsf{BV}_{q,n}$
-$\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$
+$\left|\psi\,\right\rangle\in(\mathbb{C}^{q})^{n+1}$
 2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
-3:
-4:Apply $\mathsf{QFT}_{\mathbb{Z}_{q}}^{n+1}$ on $\lvert\psi\rangle$
-5:Measure all qudits to get $({\bm{x}},y)\in\mathbb{Z}_{q}^{n}\!\times\!\mathbb{Z}_{q}$
-6:if $\gcd(y,q)=1$ then
-7:     return $\tilde{{\bm{s}}}:=-y^{-1}{\bm{x}}\bmod q$triangleright component-wise modular division
-8:end if
-9:return $\bot$
-One may think of the input $\lvert\psi\rangle$ as what we get by querying an oracle $\mathcal{O}_{\bm{s}}$. This algorithm will be the core subroutine in the learning algorithms.
-At this point, it will be a natural question to ask if the algorithm still works for noisy input samples. The main contribution of [12] is to answer the question as briefly reviewed below.
-2.3 LWE algorithm with quantum samples
-2.3.1 Quantum sample
+3:Apply $\mathsf{QFT}_{\mathbb{Z}_{q}}^{n+1}$ on $\left|\psi\,\right\rangle$
+4:Measure all qudits to get $({\bm{x}},y)\in\mathbb{Z}_{q}^{n}\!\times\!\mathbb{Z}_{q}$
+5:if $\gcd(y,q)=1$ then
+6:     return $\tilde{{\bm{s}}}:=-y^{-1}{\bm{x}}\bmod q$
+7:end if
+8:return $\bot$
+One may think of the input $\left|\psi\,\right\rangle$ as what we get by querying an oracle $\mathcal{O}_{\bm{s}}$. This algorithm will be the core subroutine in the learning algorithms.
+At this point, it will be a natural question to ask if the algorithm still works for noisy input samples.
+The main contribution of the previous work [6] is to answer the question as briefly reviewed below.
+LWE algorithm with quantum samples
 Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector and let $\mathcal{V}\subseteq\mathbb{Z}_{q}^{n}$.
-A quantum LWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$, is defined as follows:
-$\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{{\bm{a}}\in\mathcal{V}}\lvert{\bm{a}}\rangle\lvert\left<{\bm{s}},{\bm{a}}\right>+e_{\bm{a}}\bmod q\rangle,\quad e_{\bm{a}}\leftarrow\chi_{\textup{lwe}}$ (3)
-where $\chi_{\textup{lwe}}$ is a probability distribution on $\mathbb{Z}$.
-We assume that the support of $\chi_{\textup{lwe}}$ is $[-k,k]\cap\mathbb{Z}$ for some small positive integer $k$, i.e.
+A quantum LWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$, is defined as follows:
+$\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle=\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{{\bm{a}}\in\mathcal{V}}\left|{\bm{a}}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}\right>+e_{\bm{a}}\bmod q\,\right\rangle,\leavevmode\nobreak\ e_{\bm{a}}\leftarrow\chi_{\textup{\tiny lwe}}\enspace,$ (3)
+where $\chi_{\textup{\tiny lwe}}$ is a probability distribution on $\mathbb{Z}$.
+We assume that the support of $\chi_{\textup{\tiny lwe}}$ is $[-k,k]\cap\mathbb{Z}$ for some small positive integer $k$, i.e.
 $|e_{{\bm{a}}}|\leq k$.
-Define a quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ which outputs $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ when queried.
-2.3.2 Algorithm
-Given access to the quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$, one can find the secret vector ${\bm{s}}$ by using the $\mathsf{qLWE}\textup{-}\mathsf{Solver}$ described in Algorithm 2.
-Algorithm 2 $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
-1:oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
-2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
-3:
-4:$i\leftarrow 0$
-5:repeat
-6:     $\lvert\psi^{\textup{lwe}}_{{\bm{s}},V}\rangle\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright oracle query
-7:     $\tilde{{\bm{s}}}\leftarrow\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$
-8:     if  $\tilde{{\bm{s}}}=\bot$  then
-9:         return $\bot$
-10:     end if
-11:     if  $\mathsf{Test}(\tilde{{\bm{s}}},\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}})=1$ then
-12:         return $\tilde{{\bm{s}}}$
-13:     end if
-14:     $i\leftarrow i+1$
-15:until $i\leq\ell_{1}$
-16:return $\bot$
+Define a quantum oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$ which outputs $\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$ when queried.
+Given access to the quantum oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$, one can find the secret vector ${\bm{s}}$ by using the $\mathsf{qLWE}\textup{-}\mathsf{Solver}$ described in Algorithm 2 (see below).
 The subroutine $\mathsf{Test}$ in Algorithm 2 is designed to check if the candidate $\tilde{{\bm{s}}}$ is indeed the secret vector ${\bm{s}}$ by using additional LWE samples.
-Note that in Algorithm 3 the oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ can be quantum or classical.
-Algorithm 3 $\mathsf{Test}$
-1:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$, oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
-2:$1$ (pass) or 00 (fail)
-3:
-4:$i\leftarrow 0$
-5:repeat
-6:     $({\bm{a}},b)\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright $b=\left<{\bm{a}},{\bm{s}}\right>+e_{\bm{a}}\bmod q$
-7:     if $|b-\left<{\bm{a}},\tilde{{\bm{s}}}\right>\bmod q|>k$ then
-8:         return 00
-9:     end if
-10:     $i\leftarrow i+1$
-11:until $i\leq\ell_{2}$
-12:return $1$
-2.3.3 Performance
-Detailed analysis on Algorithm 2 is not covered here. Interested readers are encouraged to refer to Section IV B in [12]. Results and parameter choices are summarized below.
-Success probability of $\mathsf{BV}_{q}$
-Probability that $\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$ returns the correct secret ${\bm{s}}$ is lower bounded by
-$\mathrm{\Omega}\!\left(\frac{|\mathcal{V}|}{kq^{n}}\right)\enspace.$
-Pass/fail probability of $\mathsf{Test}$
-If $\tilde{{\bm{s}}}={\bm{s}}$, $\mathsf{Test}$ returns $1$ (pass) with certainty. If $\tilde{{\bm{s}}}\neq{\bm{s}}$, the probability that $1$ is returned is at most
-$\left(\frac{2k\!+\!1}{q}\right)^{\!\ell_{2}}\enspace.$
-Parameter choices for $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
+Note that in Algorithm 3 the oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$ can be quantum or classical.
+Detailed analysis on Algorithm 2 is not covered here.
+Interested readers are encouraged to refer to Section IV B in the previous paper [6].
+Results and parameter choices are summarized as follows.
+Probability that $\mathsf{BV}_{q,n}(\displaystyle\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle)$ returns the correct secret ${\bm{s}}$ is lower bounded by $\mathrm{\Omega}\!\left({|\mathcal{V}|}/({kq^{n}})\right)$.
+If $\tilde{{\bm{s}}}={\bm{s}}$, $\mathsf{Test}$ returns $1$ (pass) with certainty.
+If $\tilde{{\bm{s}}}\neq{\bm{s}}$, the probability that $1$ is returned is at most $\left(({2k\!+\!1})/{q}\right)^{\!\ell_{2}}$.
 By setting $\mathcal{V}=\mathbb{Z}_{q}^{n}$, $\ell_{1}=O(k\log(1/\eta))$, and $\ell_{2}=1$, we have
@@
 oracle query), $\mathsf{BV}_{q,n}$, and $\mathsf{Test}$ are assumed to be $O(1)$, $\mathsf{poly}(n,\log q)$, and $\mathsf{poly}(n,\log q)$, respectively.
-3 An RLWE algorithm with quantum samples
-In this section it will be shown that the RLWE problem with quantum samples is reduced to the LWE problem with quantum samples by efficiently transforming a RLWE sample to an LWE sample.
+Algorithm 2 $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
+1:oracle access to $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$
+2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
+3:$i\leftarrow 0$
+4:repeat
+5:     $\big{|}\psi^{\textup{\tiny lwe}}_{{\bm{s}},V}\big{\rangle}\leftarrow\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$triangleright oracle query
+6:     $\tilde{{\bm{s}}}\leftarrow\mathsf{BV}_{q,n}\big{(}\big{|}\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\big{\rangle}\big{)}$
+7:     if  $\tilde{{\bm{s}}}=\bot$  then
+8:         return $\bot$
+9:     end if
+10:     if  $\mathsf{Test}(\tilde{{\bm{s}}},\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}})=1$ then
+11:         return $\tilde{{\bm{s}}}$
+12:     end if
+13:     $i\leftarrow i+1$
+14:until $i\geq\ell_{1}$
+15:return $\bot$
+Algorithm 3 $\mathsf{Test}$
+1:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$, oracle access to $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$
+2:$1$ (pass) or 00 (fail)
+3:$i\leftarrow 0$
+4:repeat
+5:     $({\bm{a}},b)\leftarrow\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$triangleright $b=\left<{\bm{a}},{\bm{s}}\right>+e_{\bm{a}}\bmod q$
+6:     if $|b-\left<{\bm{a}},\tilde{{\bm{s}}}\right>\bmod q|>k$ then
+7:         return 00
+8:     end if
+9:     $i\leftarrow i+1$
+10:until $i\geq\ell_{2}$
+11:return $1$
+Revisiting previous works
+This section presents new results.
+The first finding is about the solvability of the RLWE problem by reducing it to LWE.
+The second and third findings involve the development of efficient classical algorithms for SIS and size-reduced LWE problems under certain assumptions.
+The purpose of the second and the third results is to show the following: if the assumptions on quantum samples taken in the previous works do hold, and then there also exist polynomial-time classical algorithms.
+A handful of classical algorithms are introduced below, but before presenting the results, we are urged to clarify the meaning of the word ‘classical’ here.
+In Result 2 and Result 3, we begin with quantum samples assumed in respective previous works [6, 8].
+We do not question or evaluate the practicality of preparing quantum samples, but rather immediately carry out the measurement in computational basis.
+Each measurement gives rise to one classical sample that is put as an input to classical algorithms we develop.
+Therefore in estimating the complexity of classical algorithms, the number of samples used should be read as the number of quantum samples assumed.
+Result 1: RLWE algorithm with quantum samples
+The goal of this subsection is to solve the RLWE problem given that the LWE problem is solvable by Algorithm 2.
+Applying Algorithm 2 to RLWE problem was originally unsuccessful by the authors of [6, Section V].
+Below the RLWE problem with quantum samples is reduced to LWE by efficiently transforming a RLWE sample to an LWE sample.
 The reduction utilizes linearity of operations in $\mathcal{R}$.
-To be more specific, a RLWE sample is transformed to an LWE sample in a way that coefficients of a polynomial becomes a vector.
-The LWE algorithm described in Section 2.3 is then applied to recover $s(X)$.
-Grilo et al.
-mentioned in their paper [12] that the LWE algorithm is not applicable to the RLWE problem due to the absence of a proper isomorphism from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$.
-Here we show that it can be done by employing the linearity.
-3.0.1 Quantum sample
+To be more specific, a RLWE sample is transformed to an LWE sample in a way that the coefficients of a polynomial become a vector.
+The LWE algorithm described in the previous section is then applied to recover $s(X)$.
 Let ${\bm{s}}(X)$ be a secret polynomial and let $\mathcal{V}\subseteq\mathcal{R}_{q}$.
-A quantum RLWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{rlwe}}_{{\bm{s}},\mathcal{V}}\rangle$ is defined as follows:
-$\lvert\psi_{s,\mathcal{V}}^{\textup{rlwe}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{a(X)\in\mathcal{V}}\lvert a(X)\rangle\lvert s(X)\!\cdot\!a(X)+e_{a}(X)\bmod q\rangle,\quad e_{a}(X)\leftarrow\chi_{\textup{rlwe}}\enspace$
-where $\chi_{\textup{rlwe}}$ is a probability distribution on $\mathcal{R}$ which samples $e_{a}(X)=\sum_{i=0}^{n-1}e_{a,i}X^{i}$ as $e_{a,0},\ldots,e_{a,n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
-3.0.2 Transform from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$
+A quantum RLWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\left|\psi^{\textup{\tiny rlwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$ is defined as follows:
+$\left|\psi_{s,\mathcal{V}}^{\textup{\tiny rlwe}}\,\right\rangle=\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{a(X)\in\mathcal{V}}\left|a(X)\,\right\rangle\left|s(X)\!\cdot\!a(X)+e_{a}(X)\bmod q\,\right\rangle,$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e_{a}(X)\leftarrow\chi_{\textup{\tiny rlwe}}\enspace,$
+where $\chi_{\textup{\tiny rlwe}}$ is a probability distribution on $\mathcal{R}$ which samples $e_{a}(X)=\sum_{i=0}^{n-1}e_{a,i}X^{i}$ as $e_{a,0},\ldots,e_{a,n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
+Transform from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$.
 Recall that the polynomial ring $\mathcal{R}$ is defined as $\mathcal{R}=\mathbb{Z}[X]/\langle\phi(X)\rangle$.
 Let us consider a (row) vector and a matrix constructed from a polynomial $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ as follows:
-$\mathsf{V}(a)=(a_{0},a_{1},\ldots,a_{n-1})\in\mathbb{Z}^{n},\quad\mathsf{M}_{\phi}(a)=\left(\begin{array}[]{ccc}&\mathsf{V}(a)&\\ &\mathsf{V}(X\!\cdot\!a\bmod\phi)&\\ &\vdots&\\ &\mathsf{V}(X^{n-1}\!\cdot\!a\bmod\phi)&\end{array}\right)\in\mathbb{Z}^{n\times n}$
+$\mathsf{V}(a)$ $=(a_{0},a_{1},\ldots,a_{n-1})\in\mathbb{Z}^{n},$
+$\mathsf{M}_{\phi}(a)$ $=\left(\begin{array}[]{ccc}&\mathsf{V}(a)&\\ &\mathsf{V}(X\cdot a\bmod\phi)&\\ &\vdots&\\ &\mathsf{V}(X^{n-1}\cdot a\bmod\phi)&\end{array}\right)\in\mathbb{Z}^{n\times n}\enspace,$
 where $\cdot$ is a multiplication of polynomials. Let $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$ and define
@@
 We then have a useful lemma.
-Lemma 1
+Lemma 1.
 For any $a(X),b(X)\in\mathcal{R}$, the following relations hold:
-$\mathsf{V}(a\cdot b\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathsf{M}_{\phi}(b)\,,$
-$\mathsf{V}(X\cdot a\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathbf{P}\,.$
-Proof
+$\mathsf{V}(a\cdot b\bmod\phi)=\mathsf{V}(a)\,\mathsf{M}_{\phi}(b)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{V}(X\cdot a\bmod\phi)=\mathsf{V}(a)\,\mathbf{P}\enspace.$
+Proof.
 Straightforwardly,
-$a(X)\cdot b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}X^{i}b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}\big{(}X^{i}b(X)\bmod\phi\big{)}$
+$a(X)\cdot b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}X^{i}b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}\big{(}X^{i}b(X)\bmod\phi\big{)}\enspace,$
 which verifies the first equation.
-Next, from $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$ we have $X^{n}\equiv\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi$, and thus
+Next, from $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$, we have $X^{n}\equiv\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi$, and thus
 $X\cdot a(X)=\sum_{i=1}^{n-1}a_{i-1}X^{i}+a_{n-1}X^{n}\equiv\sum_{i=1}^{n-1}a_{i-1}X^{i}+a_{n-1}\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi\enspace.$
 Therefore,
-$\mathsf{V}(X\cdot a\bmod\phi)$
-$=$ $(0,a_{0},\ldots,a_{n-2})+a_{n-1}(\phi_{0},\phi_{1},\ldots,\phi_{n-1})$
+$\phantom{=}\mathsf{V}(X\cdot a\bmod\phi)=(0,a_{0},\ldots,a_{n-2})+a_{n-1}(\phi_{0},\phi_{1},\ldots,\phi_{n-1})$
 $=(a_{0},\ldots,a_{n-1})\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ 0&0&0&\cdots&0\end{array}\right)+\ (a_{0},\ldots,a_{n-1})\left(\begin{array}[]{ccccc}0&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\\ \phi_{0}&\phi_{1}&\phi_{2}&\cdots&\phi_{n-1}\end{array}\right)$
-$=$ $\mathsf{V}(a)\,\mathbf{P}\enspace.$
-Before continuing, let us introduce some notations.
+$=\mathsf{V}(a)\,\mathbf{P}\enspace,$
+the second equation holds. ∎
 Let $\mathbf{M}^{T}$ denote the transpose of $\mathbf{M}\in\mathbb{R}^{l\times m}$.
@@
 which shows that
-$(\mathsf{M}_{\phi}(b)_{j})^{T}$ $=$ $\big{(}\mathsf{V}(b)_{j},(\mathsf{V}(b)\mathbf{P})_{j},\ldots,(\mathsf{V}(b)\mathbf{P}^{n-1})_{j}\big{)}$ (8)
-$=$ ($\mathsf{V}(b),((\mathbf{I}_{n})_{j})^{T}\right>,\left<\mathsf{V}(b),(\mathbf{P}_{j})^{T}\right>\ldots,\left<\mathsf{V}(b),((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\,.$
-3.0.3 RLWE to LWE
-Let ${\bm{a}}=\mathsf{V}(a)$ for $a\in\mathcal{R}$. Define a unitary operation,
-$U_{\phi,j}:\lvert{\bm{a}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle,\quad j\in\{0,\ldots,n-1\}$
-where ${\bm{a}}_{j}=(\mathsf{M}_{\phi}(a)_{j})^{T}=\big{(}\!\left<{\bm{a}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{a}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{a}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$.
+$(\mathsf{M}_{\phi}(b)_{j})^{T}=\big{(}\mathsf{V}(b)_{j},(\mathsf{V}(b)\mathbf{P})_{j},\ldots,(\mathsf{V}(b)\mathbf{P}^{n-1})_{j}\big{)}=\big{(}\!\left<\mathsf{V}(b),((\mathbf{I}_{n})_{j}\!)^{T}\!\right>,\left<\mathsf{V}(b),(\mathbf{P}_{j}\!)^{T}\!\right>\!,\ldots,\!\left<\mathsf{V}(b),((\mathbf{P}^{n-1})_{j}\!)^{T}\!\right>\!\big{)}\enspace.$ (4)
+RLWE to LWE. Let ${\bm{a}}=\mathsf{V}(a)$ for $a\in\mathcal{R}$. Define a unitary operation,
+$U_{\phi,j}:\left|{\bm{a}}\,\right\rangle\left|0\,\right\rangle\mapsto\left|{\bm{a}}\,\right\rangle\left|{\bm{a}}_{j}\,\right\rangle,\quad j\in\{0,\ldots,n-1\}\enspace,$
+where,
+${\bm{a}}_{j}=(\mathsf{M}_{\phi}(a)_{j})^{T}=\big{(}\!\left<{\bm{a}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{a}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{a}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\enspace.$
 We see that the operation is efficiently computable as any power of $\mathbf{P}$ can be classically pre-computed by knowing the fixed polynomial $\phi(X)$.
-Now let ${\bm{s}}=\mathsf{V}(s)$ and ${\bm{e}}_{a}=(e_{a,0},\ldots,e_{a,n-1})=\mathsf{V}(e_{a})$. By Lemma 1, Eq. (8), and the linearity of $\mathsf{V}$, we have
-$\mathsf{V}(s\cdot a+e_{a}\bmod\phi)$ $=$ $\mathsf{V}(s)\cdot\mathsf{M}_{\phi}(a)+\mathsf{V}(e_{a})$
-[…]
+Now let ${\bm{s}}=\mathsf{V}(s)$ and ${\bm{e}}_{a}=(e_{a,0},\ldots,e_{a,n-1})=\mathsf{V}(e_{a})$. By Lemma 1, Equation (4), and the linearity of $\mathsf{V}$, we have
+$\mathsf{V}(s\cdot a+e_{a}\bmod\phi)=\mathsf{V}(s)\cdot\mathsf{M}_{\phi}(a)+\mathsf{V}(e_{a})=\big{(}\!\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0},\left<{\bm{s}},{\bm{a}}_{1}\right>+e_{a,1},\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\big{)}\enspace.$
 For a given quantum RLWE sample, we apply the following transformations:
-$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}s(X)\!\cdot\!a(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
-$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
-$\xmapsto{U_{\phi,j}}$ $\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle\lvert\ldots,\,\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q,\,\ldots\rangle\enspace.$
-Ignoring or uncomputing irrelevant information, we have a quantum LWE sample with secret ${\bm{s}}$:
-$\lvert{\bm{a}}_{j}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q\rangle\enspace.$
-Alternatively, interchanging the role of $a(X)$ and $s(X)$ in the transformation, we have a quantum LWE sample with secret ${\bm{s}}_{j}$:
-$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}a(X)\!\cdot\!s(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
-$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{a}},{\bm{s}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{a}},{\bm{s}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
-where ${\bm{s}}_{j}:=(\mathsf{M}_{\phi}({\bm{s}})_{j})^{T}=\big{(}\!\left<{\bm{s}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{s}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{s}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$.
-4 On SIS algorithm
-In [12, Appendix B], a quantum sample for SIS is defined by,
-$\lvert\Psi^{\textup{sis}}\rangle=\frac{1}{\sqrt{q^{nm}}}\sum_{\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}}\lvert\mathbf{A}\rangle\lvert\mathbf{A}{{\bm{v}}}\bmod q\rangle$ (9)
-where $\mathbf{A}=[\,{\bm{a}}_{0}\,|\,\cdots\,|\,{\bm{a}}_{n-1}\,]$, ${\bm{a}}_{i}\in\mathbb{Z}_{q}^{m\times 1}$, and ${\bm{v}}\in\mathbb{Z}^{n}$ is a short secret vector such that $0<\|{\bm{v}}\|\leq\beta$.
+$\left|a(X)\,\right\rangle\left|0\,\right\rangle\left|s(X)\!\cdot\!a(X)+e_{a}(X)\bmod\phi\,\right\rangle$ $\stackrel{{\scriptstyle\mathsf{V}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{0}}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0},\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\,\right\rangle$
+$\stackrel{{\scriptstyle U_{\phi,j}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{a}}_{j}\,\right\rangle\left|\ldots,\,\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j},\,\ldots\,\right\rangle\enspace,$
+where $\bmod\;q$ is omitted in the third register. Uncomputing irrelevant information, we have a quantum sample with secret ${\bm{s}}$:
+$\left|{\bm{a}}_{j}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q\,\right\rangle\enspace.$
+Alternatively, interchanging the role of $a(X)$ and $s(X)$ in the transformation, we have a quantum sample with secret ${\bm{s}}_{j}$:
+$\left|a(X)\,\right\rangle\left|0\,\right\rangle\left|a(X)\!\cdot\!s(X)+e_{a}(X)\bmod\phi\,\right\rangle\stackrel{{\scriptstyle\mathsf{V}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{0}}\,\right\rangle\left|\left<{\bm{a}},{\bm{s}}_{0}\right>+e_{a,0},\ldots,\left<{\bm{a}},{\bm{s}}_{n-1}\right>+e_{a,n-1}\,\right\rangle\enspace,$
+where,
+${\bm{s}}_{j}=(\mathsf{M}_{\phi}({\bm{s}})_{j})^{T}=\big{(}\!\left<{\bm{s}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{s}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{s}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\enspace.$
+In either case, the LWE algorithm is then applied to solve the problem.
+Result 2: A classical algorithm for SIS problem
+We develop an efficient classical algorithm for solving SIS given that the following quantum sample is provided. In [6, Appendix B], a quantum sample for SIS is defined by,
+$\left|\Psi^{\textup{\tiny sis}}\,\right\rangle=\frac{1}{\sqrt{q^{nm}}}\sum_{\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}}\left|\mathbf{A}\,\right\rangle\left|\mathbf{A}{{\bm{v}}}\bmod q\,\right\rangle\enspace,$ (5)
+where ${\bm{v}}\in\mathbb{Z}^{n}$ is a short secret vector such that $0<\|{\bm{v}}\|\leq\beta$.
 The authors have developed a polynomial time quantum algorithm for finding ${\bm{v}}$ by using $O(2n\beta)$ quantum samples.
@@
 Instead, let us focus on the problem itself.
-From the definition of the SIS problem in Section 2.1.4, we are given an under-determined system of equations ($m<n$).
-Hardness of the SIS problem comes from the fact that among the exponentially many solutions ($q^{n-m}$), there exists no known way to choose a short one within polynomial time.
-Having ability to prepare the quantum sample multiple times for the same ${\bm{v}}$ thus already means their is a way to choose the short solution out of many solutions.
-On the other hand, the hardness of LWE comes from difficulty of solving noisy linear systems.
-Distinct LWE algorithms may require different number of noisy equations, but accessing more than one noisy equation is not implausible.
-Assuming the SIS quantum sample in Eq.
-(9) is obtainable, a classical algorithm can be designed.
-By classical, we mean no quantum process is required except measurements.
-Let $t=O(n/m)$, and for given $t$ samples just measure the state to have
+From the definition of the SIS problem, we are given an under-determined system of equations ($m<n$).
+Hardness of the SIS problem comes from the fact that among the exponentially many solutions ($\Omega(q^{n-m})$), there exists no known way to choose a short one within polynomial time.
+Assuming the SIS quantum sample in Equation (5) is obtainable, an efficient classical algorithm can be designed.
+Notice that by measuring a quantum sample, one obtains a classical sample $(\mathbf{A},\mathbf{A}{{\bm{v}}})\in\mathbb{Z}_{q}^{m\times n}\times\mathbb{Z}_{q}^{m}$.
+Prepare $t=\Omega(n/m)$ classical samples of the form $({\bm{A}}^{(i)},{\bm{z}}^{(i)})\in\mathbb{Z}_{q}^{m\times n}\times\mathbb{Z}_{q}^{m}$ which satisfies ${\bm{A}}^{(i)}{\bm{v}}\equiv{\bm{z}}^{(i)}\bmod q$ for some nonzero vector ${\bm{v}}\in\mathbb{Z}^{n}$ with $\|{\bm{v}}\|\leq\beta$.
+Each classical sample leads to $m$ linear equations (be reminded that one quantum sample leads to one classical sample, and one classical sample here corresponds to $m$ linear equations),
+$A^{(i)}_{11}v_{1}+\cdots+A^{(i)}_{1n}v_{n}\equiv z^{(i)}_{1}\bmod q,\hskip 2.84526pt\ldots\hskip 2.84526pt,\,A^{(i)}_{m1}v_{1}+\cdots+A^{(i)}_{mn}v_{n}\equiv z^{(i)}_{m}\bmod q,$
+where $A^{(i)}_{kl}$ is the $k$th row, $l$th column element of $\mathbf{A}^{(i)}$, and $v_{k}$, $z_{k}^{(i)}$ are the $k$th elements of ${\bm{v}}$ and ${\bm{z}}^{(i)}$, respectively.
+Therefore the total number of linear equations reads $mt=\Omega(n)$.
+Since there are $n$ unknowns while we have $\Omega(n)$ equations, we can find the secret vector ${\bm{v}}$ by solving the system of equations using Gaussian elimination within $O(n^{3})$ time.
+The total number of (quantum) samples used is $\Theta(n/m)$.
+Result 3: Classical algorithms for size-reduced LWE problem
+Song et al.
+have published a paper in 2022 which improved the result of Grilo et al.
+assuming a resized sample [8].
+In a nutshell, the improved algorithm and Algorithm 2 work similarly except for the resized samples.
+Therefore, here we drop the details on the algorithm and jump right into the discussion on the resized sample.
+Recall first that an LWE sample is $({\bm{a}},b=\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$, $|e|\leq k$.
+Let $\mathcal{V}_{j}\subseteq\mathbb{Z}_{q}$, ${\bm{s}}=(s_{0},\ldots,s_{n-1})\in\mathbb{Z}_{q}^{n}$, and consider the noise model $\chi_{\textup{\tiny lwe}}$ as a discrete uniform or a bounded Gaussian distribution in the interval $[-\xi,\xi]$ around zero with $\xi\ll q$.
+Assume a resized quantum sample examined by [8, Equation (9)] is given,
+$\big{|}\psi_{\mathcal{V}_{j}}\big{\rangle}=\frac{1}{\sqrt{|\mathcal{V}_{j}|}}\sum_{a_{j}^{\prime}\in\mathcal{V}_{j}}\left|a_{j}^{\prime}\,\right\rangle\left|a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q\,\right\rangle,$ (6)
+where $e_{j}^{\prime}$ is derived according to [8, Equation (4)-(8)] satisfying,
+$|e_{j}^{\prime}|\leq\xi^{\prime}=O(n^{3}\xi)\ll q\enspace.$
+Below two classical algorithms are introduced assuming the resized samples are given.
+Search on the error
+We first introduce an algorithm that exhaustively searches the ‘correct error’.
+Given a resized sample, an immediate (classical) strategy would be to exhaustively search $s_{j}$ with $O(qn)$ complexity (in which the scenario is indeed discussed in [8, Discussion]), but a more sophisticated search would be to examine the error.
+Consider an oracle $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$ which on query returns the (measured) resized sample of the $j$th component $(a_{j}^{\prime},b_{j}^{\prime})$, where $b_{j}^{\prime}=a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$.
+Given access to the oracle, Algorithm 4 is developed.
+Algorithm 4 $\mathsf{SearchError}$
+1:error bound $\xi^{\prime}$, oracle access to $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
+2:$\tilde{s}_{j}\in\mathbb{Z}_{q}$
+3:$(a_{j}^{\prime},b_{j}^{\prime})\leftarrow\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
+4:${\rm e}\leftarrow-\xi^{\prime}$
+5:repeat
+6:     $\tilde{e}_{j}^{\prime}\leftarrow\rm{e}$
+7:     $\tilde{s}_{j}\leftarrow(a_{j}^{\prime})^{-1}(b_{j}^{\prime}-\tilde{e}_{j}^{\prime})\bmod q$
+8:     if  $\mathsf{e\textup{-\!}Test}\big{(}\tilde{s}_{j},\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}\big{)}=1$ then
+9:         return $\tilde{s}_{j}$
+10:     end if
+11:     $\rm{e}\leftarrow\rm{e}+1$
+12:until ${\rm e}>\xi^{\prime}$
+13:return $1$
+Algorithm 5 $\mathsf{e\textup{-\!}Test}$
+1:$\tilde{s}_{j}\in\mathbb{Z}_{q}$, oracle access to $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
+2:$1$ (pass) or 00 (fail)
+3:$i\leftarrow 0$
+4:repeat
+5:     $(a_{j}^{\prime\prime},b_{j}^{\prime\prime})\leftarrow\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
+6:     if $|b_{j}^{\prime\prime}-a_{j}^{\prime\prime}\tilde{s}_{j}\bmod q|>\xi^{\prime}$ then
+7:         return 00
+8:     end if
+9:     $i\leftarrow i+1$
+10:until $i\geq M$
+11:return $1$
+If $\tilde{e}_{j}^{\prime}=e_{j}^{\prime}$, then replacing $b_{j}^{\prime}$ in Line 5 by $a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$ affirms $\tilde{s}_{j}=s_{j}$.
+Since the error is exhaustively searched within the range $-\xi^{\prime}\leq e_{j}^{\prime}\leq\xi^{\prime}$, Algorithm 4 eventually outputs the correct $\tilde{s}_{j}(=s_{j})$ with time complexity $O(\xi^{\prime})$, as long as the probability that wrong $\tilde{s}_{j}(\neq s_{j})$ passes the test is low enough.
+If $\tilde{s}_{j}\neq s_{j}$, for a random sample $(a_{j}^{\prime\prime},b_{j}^{\prime\prime})$, the probability that $|b_{j}^{\prime\prime}-a_{j}^{\prime\prime}\tilde{s}_{j}\bmod q|\leq\xi^{\prime}$ holds is $(2\xi^{\prime}+1)/q$.
+Therefore, such $\tilde{s}_{j}\neq s_{j}$ will pass $\mathsf{e\textrm{-}Test}$ with the probability $((2\xi^{\prime}+1)/q)^{M}$.
+In the previous work [8], the probability that the BV subroutine outputs $\tilde{s}_{j}=s_{j}$ is $\Omega(1/\xi^{\prime})$ ([8, Equation (17)] with $|v_{j}|\approx q$), thus $O(\xi^{\prime})$ repetitions of the subroutine would give the correct $\tilde{s}_{j}(=s_{j})$.
+The probability that their $M$-trial test accepts an incorrect answer is at most $((2\xi^{\prime}+1)/q)^{M}$ as in [8, Equation (12)].
+One may notice that the two algorithms have similar upper bounds on the number of repetitions (of error guessing in Algorithm  4 and BV subroutine in  [8, Section 3]).
+Moreover, the probability that an incorrect candidate passes the $\mathsf{e\textrm{-}Test}$ and $M$-trial test is exactly the same.
+Direct comparison of the two algorithms is not immediate due to the differences in approaches (e.g., exhaustive or else) and unit computations (quantum or classical), but it is plausible to assess the two as similar in performance.
+Except for the query, Algorithm 4 works classically.
+Lattice reduction
+There is another way to find $s_{j}^{\prime}$ using techniques in lattice theory.
+Given two classical samples $(a_{j,k}^{\prime},b_{j,k}^{\prime}=a_{j,k}^{\prime}s_{j}+e_{j,k}^{\prime}\bmod q)$ for $k=1,2$, consider two-dimensional vectors ${\bm{a}}_{j}^{\prime}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$, ${\bm{b}}_{j}^{\prime}=(b_{j,1}^{\prime},b_{j,2}^{\prime})$, ${\bm{e}}_{j}^{\prime}=(e_{j,1}^{\prime},e_{j,2}^{\prime})$.
+Then by the construction, we know there is a vector ${\bm{v}}\in\mathbb{Z}^{2}$ such that ${\bm{v}}\equiv s_{j}{\bm{a}}_{j}^{\prime}\bmod q$ and $\|{\bm{b}}_{j}^{\prime}-{\bm{v}}\|=\|{\bm{e}}_{j}^{\prime}\|$.
+In particular, the vector ${\bm{v}}$ is very close to ${\bm{b}}_{j}^{\prime}$ since $\|{\bm{e}}_{j}^{\prime}\|\leq\sqrt{2}\xi^{\prime}$ is small.
+Furthermore, notice that a set $\{{\bm{w}}\in\mathbb{Z}^{2}:{\bm{w}}\equiv x{\bm{a}}_{j}^{\prime}\bmod q,x\in\mathbb{Z}\}$ constitutes a lattice in which ${\bm{v}}$ is included.
+In other words, for given ${\bm{b}}_{j}^{\prime}$, finding $s_{j}$ can be solved by looking for a lattice vector that is closest to ${\bm{b}}_{j}^{\prime}$, known as the closest vector problem (CVP).
+In general, the CVP is hard to solve, that is, it takes time exponential in the lattice dimension.
+However in our case, it is sufficient to solve the CVP over a two-dimensional lattice, which is tractable by using the Babai’s nearest plane algorithm with a short basis computed by the Lagrange-Gauss algorithm.
+Lagrange-Gauss algorithm.
+This algorithm is useful for finding a shortest basis of a lattice in two-dimension.
+We say that an ordered basis ${\bm{g}}_{1},{\bm{g}}_{2}$ of a lattice $L$ is Lagrange-Gauss reduced if $\|{\bm{g}}_{1}\|\leq\|{\bm{g}}_{2}\|\leq\|{\bm{g}}_{2}+z{\bm{g}}_{1}\|$ for all $z\in\mathbb{Z}$.
+It can be shown that if a basis ${\bm{g}}_{1},{\bm{g}}_{2}$ of a lattice $L$ is Lagrange-Gauss reduced, then $\|{\bm{g}}_{i}\|=\lambda_{i}(L)$, where $\lambda_{i}(L)$ is the $i$th successive minimum of $L$ [20, Chapter 17].
+This basis can be computed efficiently by Algorithm 6.
+Algorithm 6 $\mathsf{LagrangeGauss}$
+1:basis ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathbb{R}^{2}$ for a lattice $L$ such that $\|{\bm{g}}_{1}\|<\|{\bm{g}}_{2}\|$
+2:basis ${\bm{g}}_{1},{\bm{g}}_{2}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(L)$
+3:while $\|{\bm{g}}_{1}\|<\|{\bm{g}}_{2}\|$ do
+4:     ${\bm{g}}_{2}\leftarrow{\bm{g}}_{2}-\left\lfloor\!\frac{\langle{\bm{g}}_{1},{\bm{g}}_{2}\rangle}{\|{\bm{g}}_{1}\|^{2}}\!\right\rceil{\bm{g}}_{1}$
+5:     Swap ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$
+6:end while
+7:return ${\bm{g}}_{1},{\bm{g}}_{2}$
+Detailed analysis of the algorithm can be found in [20, Chapter 17]. The time complexity of the algorithm is known to be $O(\log^{3}(\max_{i}{\|{\bm{g}}_{i}\|^{2}}))$, where ${\bm{g}}_{i}$ are inputs.
+Nearest plane algorithm.
+The algorithm is for solving CVP.
+Let ${\bm{g}}_{1},\ldots,{\bm{g}}_{n}\in\mathbb{R}^{n}$ be basis vectors of a lattice $L$.
+Given a target vector ${\bm{t}}\in\mathbb{R}^{n}$, we are asked to find a vector ${\bm{v}}\in L$ such that $\|{\bm{t}}-{\bm{v}}\|$ is minimized.
+Babai’s nearest plane algorithm gives an approximate solution to CVP [21] in general dimension.
+Intuitively, what it does is to fix some basis vectors and then find a linear combination of remaining vectors such that it is closest to the projected ${\bm{u}}$ onto the hyperplane spanned by the fixed basis.
+Algorithm 7 describes the nearest plane algorithm in dimension $2$.
+We refer to [20, Chapter 18] for general dimension.
+Algorithm 7 $\mathsf{NearestPlane}$
+1:basis ${\bm{g}}_{1},{\bm{g}}_{2}$ for a lattice $L\subseteq\mathbb{R}^{2}$, target ${\bm{t}}\in\mathbb{R}^{2}$
+2:${\bm{v}}\in L$
+3:$\mathbf{g}_{1}^{\ast}\leftarrow{\bm{g}}_{1}$
+4:$\mathbf{g}_{2}^{\ast}\leftarrow{\bm{g}}_{2}-\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\mathbf{g}_{1}^{\ast}$
+5:${\bm{t}}_{2}\leftarrow{\bm{t}}$
 […]
-Notice that the total number of linear equations we have is $mt$, which is greater than $n$.
-With high probability this system of equations has the unique solution ${\bm{v}}$ in which Gaussian elimination can be applied to solve.
-5 On the Divide-and-Conquer algorithm
-Song et al.
-have published a paper in 2022 which improves the result of [12] assuming a resized sample [19].
-In a nutshell, the improved algorithm and one in Section 2.3 work similarly except the samples.
-Therefore, here we drop the details on the algorithm and jump right into the discussion on Eq.
-(9) in [19], which introduces the resized sample.
-Recall first that an LWE sample is $({\bm{a}},b=\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$, $|e|\leq k$.
-5.1 Resized sample assumption
-Let $\mathcal{V}_{j}\subseteq\mathbb{Z}_{q}$, ${\bm{s}}=(s_{0},\ldots,s_{n-1})\in\mathbb{Z}_{q}^{n}$, and assume the following sample is obtainable,
-$\lvert\psi_{\mathcal{V}_{j}}\rangle=\frac{1}{\sqrt{|\mathcal{V}_{j}|}}\sum_{a_{j}^{\prime}\in\mathcal{V}_{j}}\lvert a_{j}^{\prime}\rangle\lvert a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q\rangle,\quad|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)\ll q$ (10)
-where $e_{j}^{\prime}$ can be considered as an error from some distribution originated from the original LWE sample.
-Apparently some more assumptions are required (including QRAM) for the above statement to hold, but let us take all of them positively and Eq.
-(10) is indeed prepared.
-[^2] To us, the description of how this kind of sample is obtained in [19] is not transparent.
-In the following subsections, two classical algorithms for LWE problem is developed assuming the resized sample is given.
-5.2 Search on the Error
-Given a resized sample, a naive classical strategy is to exhaustively search $s_{j}$ with $O(qn)$ complexity (in which the scenario is indeed discussed in [19]), but more sophisticated search would be to examine the error.
-The (measured) resized sample of the $j$th component reads $b_{j}^{\prime}=a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$, $|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)$.
-Let us consider the following procedure:
-  1. 1.
-    Set a value of $e_{j}^{\prime}$ such that $|e_{j}^{\prime}|\leq k^{\prime}$.
-  2. 2.
-    For a measured sample, compute $\tilde{s}_{j}:=(a_{j}^{\prime})^{-1}(b_{j}^{\prime}-e_{j}^{\prime})\bmod q$.
-  3. 3.
-    For $1\leq i\leq M$ for some integer $M$, test if $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$, where $b_{j,i}^{\prime}$ and $a_{j,i}^{\prime}$ are $i$th sample for the test on the $j$th component.
-  4. 4.
-    If it passes the test, return $\tilde{s}_{j}$. Otherwise, set a different value for $e_{j}^{\prime}$ and repeat.
-Let us estimate the complexity.
-If $\tilde{s}_{j}=s_{j}$, then $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is always true and it will pass.
-The probability that a candidate $e_{j}^{\prime}$ leads to the correct $\tilde{s}_{j}$ is thus at least $\Omega(1/k^{\prime})$.
-On the other hand if $\tilde{s}_{j}\neq s_{j}$, then the probability that $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is true is $(2k^{\prime}+1)/q$, thus $\tilde{s}_{j}\neq s_{j}$ will pass with the probability $((2k^{\prime}+1)/q)^{M}$.
-Compare it with the algorithm in [19].
-In [19], the probability that the BV subroutine outputs $\tilde{s}_{j}=s_{j}$ is $\Omega(1/k^{\prime})$ (Eq.
-(17) in [19] with $|v_{j}|\approx q$).
-The probability that $M$-trial test accepts an incorrect answer is at most $((2k^{\prime}+1)/q)^{M}$ as in Eq.
-(12) in [19].
-It is not difficult to see that the above classical procedure is at least as efficient as the algorithm proposed in [19].
-5.3 Lattice reduction
-Given two (measured) resized samples $(a_{j,1}^{\prime},b_{j,1}^{\prime}=a_{j,1}^{\prime}s_{j}+e_{j,1}^{\prime}\bmod q)$, $(a_{j,2}^{\prime},b_{j,2}^{\prime}=a_{j,2}^{\prime}s_{j}+e_{j,2}^{\prime}\bmod q)$, let us consider the following lattice:
-$\mathcal{L}$ $=$ $\big{\{}{\bm{v}}\in\mathbb{Z}^{2}:{\bm{v}}\equiv x{\bm{a}}\bmod q\textup{ for some }x\in\mathbb{Z}\big{\}}$
-where ${\bm{a}}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$ and ‘$\bmod~{}q$’ applies to each component.
-Assume that $a_{j,1}^{\prime}$ is invertible in $\mathbb{Z}_{q}$, then $(1,(a_{j,1}^{\prime})^{-1}a_{j,2}\bmod q),(0,q)$ is a basis of $\mathcal{L}$ and so $\dim(\mathcal{L})=2$, $\det(\mathcal{L})=q$.
-The purpose of introducing $\mathcal{L}$ is to reduce finding $s_{j}$ to the closest vector problem (CVP).
-Notice that if ${\bm{e}}:=(e_{j,1}^{\prime},e_{j,2}^{\prime})={\bm{0}}$, then ${\bm{b}}:=(b_{j,1}^{\prime},b_{j,2}^{\prime})\in\mathcal{L}$.
-For nonzero small $\|{\bm{e}}\|$, we thus expect that ${\bm{b}}\equiv s_{j}{\bm{a}}+{\bm{e}}\bmod q$ is close to a vector $s_{j}{\bm{a}}\bmod q\in\mathcal{L}$.
-Therefore the task is to solve CVP in $\mathcal{L}$ for target ${\bm{b}}$.
-Generally the CVP is taken to be a hard problem, but in two dimension there exists an efficient solution: Lagrange-Gauss reduction is first applied to find a good basis of $\mathcal{L}$, and then Babai’s nearest plane algorithm is used to find the closest vector in $\mathcal{L}$ to ${\bm{b}}$ [3].
-By applying the Lagrange-Gauss reduction algorithm, one can compute basis vectors ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathcal{L}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(\mathcal{L})$ in time $O(\log^{3}q)$ [11, Chapter 17], where $\lambda_{i}(\mathcal{L})$ is the $i$-th successive minimum of $\mathcal{L}$.
-For given basis vectors $\mathbf{b}_{1},\ldots,\mathbf{b}_{\dim(\mathcal{L})}$ and a target ${\bm{u}}$, if there is a vector ${\bm{v}}\in\mathcal{L}$ satisfying $\|{\bm{v}}-{\bm{u}}\|\leq\frac{1}{2}\min_{i\in\{1,\ldots,\dim(\mathcal{L})\}}\{\|\mathbf{b}_{i}^{*}\|\}$, then the Babai’s nearest plane algorithm always outputs the vector ${\bm{v}}$ in time $O(\log^{2}q)$ [11, Chapter 18], where $\mathbf{b}_{i}^{*}$ is the $i$-th Gram-Schmidt orthogonalized vector.
-Assuming the Gaussian heuristic [11, Chapter 16] holds, $\|{\bm{g}}_{1}\|\approx\sqrt{\det(\mathcal{L})}=\sqrt{q}$.
-Since ${\bm{g}}_{1}={\bm{g}}_{1}^{*}$ and $\|{\bm{g}}_{1}^{*}\|\|{\bm{g}}_{2}^{*}\|=\det(\mathcal{L})$, we have $\|{\bm{g}}_{2}^{*}\|\approx\sqrt{q}$.
-Therefore Babai’s algorithm results in a vector ${\bm{v}}\in\mathcal{L}$ such that $\|{\bm{v}}-{\bm{b}}\|\leq\frac{1}{2}\sqrt{q}$.
-If $\|{\bm{e}}\|\leq\frac{1}{2}\sqrt{q}$, then one would expect ${\bm{v}}\equiv s_{j}{\bm{a}}\bmod q$.
-Indeed, since $|e_{j,i}^{\prime}|\leq k^{\prime}$ (i.e., $\|{\bm{e}}\|\leq\sqrt{2}k^{\prime}$), a condition for the above strategy to work is
-$8k^{\prime 2}\leq q\enspace.$ (11)
-Equation (6) in [19] assumes $k^{\prime}\ll q$, thus Eq. (11) likely hold.
+7:${\bm{t}}_{1}\leftarrow{\bm{t}}_{2}-z_{2}{\bm{g}}_{2}$
+[…]
+9:return ${\bm{v}}=z_{1}{\bm{g}}_{1}+z_{2}{\bm{g}}_{2}$
+Proposition 1.
+Let ${\bm{g}}_{1},{\bm{g}}_{2}\in L$, and ${\bm{t}}\in\mathbb{R}^{2}$ be inputs to Algorithm 7 and let $\mathbf{g}_{1}^{*},\mathbf{g}_{2}^{*}$ be Gram-Schmidt orthogonalized vectors of ${\bm{g}}_{1},{\bm{g}}_{2}$.
+If there exists ${\bm{w}}\in L$ such that $\|{\bm{t}}-{\bm{w}}\|<\frac{1}{2}\min_{i\in\{1,2\}}(\|\mathbf{g}_{i}^{\ast}\|)$, then Algorithm  7 outputs ${\bm{w}}$.
+Proof.
+Let ${\bm{v}}=z_{1}{\bm{g}}_{1}+z_{2}{\bm{g}}_{2}$ be the output of Algorithm 7, and let ${\bm{t}}_{1},{\bm{t}}_{2}$ be vectors as in the algorithm.
+Assume there exists ${\bm{w}}$ such that ${\bm{w}}=w_{1}{\bm{g}}_{1}+w_{2}{\bm{g}}_{2}={\bm{t}}-{\bm{e}}$ for some $w_{1},w_{2}\in\mathbb{Z}$ and $\|{\bm{e}}\|<\frac{1}{2}\min_{i\in\{1,2\}}(\|\mathbf{g}_{i}^{\ast}\|)$.
+Then we have a useful bound on $|\langle{\bm{e}},\mathbf{g}_{i}^{\ast}\rangle|/\|\mathbf{g}_{i}^{\ast}\|^{2}$ for $i=1,2$,
+$\frac{|\langle{\bm{e}},\mathbf{g}_{i}^{\ast}\rangle|}{\|\mathbf{g}_{i}^{\ast}\|^{2}}\leq\frac{\|{\bm{e}}\|\|\mathbf{g}_{i}^{\ast}\|}{\|\mathbf{g}_{i}^{\ast}\|^{2}}<\frac{1}{2}\enspace.$ (7)
+Recall that ${\bm{t}}_{2}={\bm{t}}={\bm{w}}+{\bm{e}}$. From
+$\frac{\langle{\bm{t}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=\frac{\langle{\bm{w}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=\frac{\langle w_{1}{\bm{g}}_{1}+w_{2}{\bm{g}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=w_{2}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}\enspace,$
+we see that $z_{2}=\left\lfloor\frac{\langle{\bm{t}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}\right\rceil=w_{2}$ due to Equation (7).
+Similarly, recalling ${\bm{t}}_{1}={\bm{t}}-z_{2}{\bm{g}}_{2}$ and from
+$\frac{\langle{\bm{t}}-z_{2}{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}$ $=\frac{\langle{\bm{w}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}-z_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}$
+$=w_{1}+w_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}-z_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}=w_{1}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\enspace,$
+we have $z_{1}=\left\lfloor\frac{\langle{\bm{t}}_{1},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\right\rceil=w_{1}$, proving ${\bm{v}}={\bm{w}}$. ∎
+The time complexity of Algorithm 7 is $O(\log^{2}B)$ where $B=\max\{\|{\bm{g}}_{1}\|,\|{\bm{g}}_{2}\|\}$.
+Finding $s_{j}$ by lattice reduction.
+In addition to the above algorithms, we adopt Gaussian heuristic which predicts that the number of lattice points inside a measurable set $\mathcal{B}\subset\mathbb{R}^{n}$ is approximately equal to $\operatorname{Vol}(\mathcal{B})/\operatorname{Vol}(L)$ where $L$ is a full-rank lattice in $\mathbb{R}^{n}$ and $\operatorname{Vol}(L):=\det(L)$.
+Gaussian heuristic is widely used in analyzing the performance of lattice-related algorithms.
+Applied to Euclidean ball in dimension $n$, we expect the length of a nonzero shortest vector in a random lattice $L$ to be approximately,
+$\sqrt{\frac{n}{2\pi e}}\det(L)^{1/n}\enspace.$
+We are now ready to describe the method.
+Suppose we are given two classical samples $(a_{j,k}^{\prime},b_{j,k}^{\prime}=a_{j,k}^{\prime}s_{j}+e_{j,k}^{\prime}\bmod q)$ for $k=1,2$, by measuring the resized quantum samples.
+Assume that $a_{j,1}^{\prime}$ is invertible in $\mathbb{Z}_{q}$.
+Let ${\bm{a}}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$ and consider the following lattice:
+=$\mathcal{L}=\big{\{}{\bm{v}}\in\mathbb{Z}^{2}:{\bm{v}}\equiv x{\bm{a}}\bmod q\textup{ for some }x\in\mathbb{Z}\big{\}}\enspace,$
+where $\bmod\leavevmode\nobreak\ q$ applies to each component.
+Notice that by letting ${\bm{v}}=(v_{1},v_{2})\equiv x{\bm{a}}$ for some $x\in\mathbb{Z}$, we have $x\equiv(a_{j,1}^{\prime})^{-1}v_{1}\bmod q$.
+This shows that two vectors $\mathbf{b}_{1}=(1,(a_{j,1}^{\prime})^{-1}a_{j,2}^{\prime}\bmod q)$ and $\mathbf{b}_{2}=(0,q)$ form a basis of $\mathcal{L}$, that is,
+=$\mathcal{L}=\{{\bm{v}}:{\bm{v}}=z_{1}\mathbf{b}_{1}+z_{2}\mathbf{b}_{2},\leavevmode\nobreak\ z_{1},z_{2}\in\mathbb{Z}\},$
+and thus $\dim(\mathcal{L})=2$, $\det(\mathcal{L})=\det\!\big{(}\big{[}{\mathbf{b}_{1}\atop\mathbf{b}_{2}}\big{]}\big{)}=q$.
+As explained earlier, the purpose of introducing $\mathcal{L}$ is to reduce finding $s_{j}$ to solving CVP: if one manages to find ${\bm{v}}\in\mathcal{L}$ that is close to ${\bm{b}}=(b_{j,1}^{\prime},b_{j,2}^{\prime})$, then the error is obtained by ${\bm{e}}={\bm{b}}-{\bm{v}}$.
+By using Lagrange-Gauss algorithm with inputs $\mathbf{b}_{1},\mathbf{b}_{2}$, we have basis vectors ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathcal{L}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(\mathcal{L})$ for $i=1,2$.
+Assuming the Gaussian heuristic holds, $\lambda_{1}(\mathcal{L})=\|{\bm{g}}_{1}\|\approx\sqrt{\det(\mathcal{L})}=\sqrt{q}$.
+Let $\mathbf{g}_{1}^{*},\mathbf{g}_{2}^{*}$ be Gram-Schmidt orthogonalized vectors of ${\bm{g}}_{1},{\bm{g}}_{2}$.
+Due to the properties of Gram-Schmidt orthogonalized vectors, we know that ${\bm{g}}_{1}=\mathbf{g}_{1}^{*}$ and $\det(\mathcal{L})=\det\!\big{(}\big{[}{\mathbf{b}_{1}\atop\mathbf{b}_{2}}\big{]}\big{)}=\det\!\big{(}\big{[}{\mathbf{g}_{1}^{\ast}\atop\mathbf{g}_{2}^{\ast}}\big{]}\big{)}=\|\mathbf{g}_{1}^{*}\|\|\mathbf{g}_{2}^{*}\|$.
+Therefore, $\|\mathbf{g}_{2}^{*}\|\approx\sqrt{q}$.
+Since ${\bm{b}}\equiv s_{j}{\bm{a}}+{\bm{e}}\bmod q$, we have ${\bm{b}}=s_{j}{\bm{a}}+{\bm{e}}+q{\bm{l}}$ for some ${\bm{l}}\in\mathbb{Z}^{2}$ and $\|{\bm{b}}-(s_{j}{\bm{a}}+q{\bm{l}})\|=\|{\bm{e}}\|$.
+Thus, if $\|{\bm{e}}\|<\frac{1}{2}\min\{\|\mathbf{g}_{1}^{\ast}\|,\|\mathbf{g}_{2}^{\ast}\|\}\approx\frac{1}{2}\sqrt{q}$, then taking ${\bm{g}}_{1},{\bm{g}}_{2}$ and ${\bm{b}}$ as inputs, Babai’s algorithm outputs the vector ${\bm{w}}=s_{j}{\bm{a}}+q{\bm{l}}$ by Proposition 1.
+Computing ${\bm{w}}\bmod q$ then yields $s_{j}{\bm{a}}\bmod q$, from which we recover $s_{j}$.
+Indeed, since $|e_{j,i}^{\prime}|\leq\xi^{\prime}$ (i.e., $\|{\bm{e}}\|\leq\sqrt{2}\xi^{\prime}$), a condition for the algorithm to work is
+$8\xi^{\prime 2}\lesssim q\enspace.$ (8)
+In the previous paper, the authors assume $\xi^{\prime}\ll q$, thus Equation (8) likely holds.
+Considering the complexity of each subroutine algorithm, the total running time is $\mathsf{poly}(\log(q))$.
+Figure 2: Success rate of the proposed algorithm as a function of the error bound for (a) $q=3329$ (Kyber) and for (b) $q=2^{23}-2^{13}+1$ (Dilithium).
+In each figure, the condition given by Equation (8) is around 20 and 1023, respectively.
+In (b), the interval in the horizontal axis is 10.
+The above analysis relies on Gaussian heuristic which is not guaranteed to work beneficially on every instance.
+To verify the validity of the approach, numerical tests are carried out summarized in Figure 2.
+Each subfigure shows the success rate of the lattice reduction method for chosen modulus $q$ from Kyber [22] and Dilithium [23], respectively.
+For each error bound $\xi^{\prime}$, the following procedure is conducted 1000 times: a pair of resized samples is produced randomly (including the secret) and the algorithm is applied to output a candidate for the secret.
+The success rate is the number of times the algorithm correctly outputs the secret divided by 1000.
+As can be seen from the figure, the algorithm succeeds better as the bound gets smaller.
+There is a chance that the algorithm fails for small $\|{\bm{e}}\|$ due to unusually short Gram-Schmidt vectors, but such cases can be overcome by getting another sample and trying a different lattice.
+Conclusion
+A definitive conclusion on whether the linear learning algorithms will be immediately useful once a powerful quantum computer is ready is not yet drawn from this work, but we have shown that the RLWE problem can be tackled if LWE is indeed solvable quantumly.
+In addition, it is also shown that assumptions taken in solving SIS and size-reduced LWE problems lead to efficient classical algorithms, implying that the previously claimed quantum advantages need to be reexamined.
+A further multidisciplinary study on the subject may estimate the true feasibility of the algorithms.
+Data availability
+All data generated or analysed during this study are included in this published article.